### v3

# On robustness of Spectral Rényi divergence

Abstract
This paper studies a specific category of statistical divergences for spectral densities of time series: the spectral $\alpha$-Rényi divergences, which includes the Itakura–Saito divergence as a subset. The aim of this paper is to highlight both information-theoretic and statistical properties of spectral $\alpha$-Rényi divergences. We reveal the connection between the spectral $\alpha$-Rényi divergence and the $\gamma$-divergence in robust statistics, and a variational representation of spectral $\alpha$-Rényi divergence. Inspired by these results suggesting “robustness” of spectral $\alpha$-Rényi divergence, we show that the minimum spectral Rényi divergence estimate has a stable optimization path with respect to outliers in the frequency domain, unlike the minimum Itakura–Saito divergence estimator, and thus it delivers more stable estimate, reducing the need for intricate pre-processing.
keywords:
Frequency domain analysis; Optimization theory; Robust statistics; Spectral density; Statistical divergence; Time series analysis
and
1 Introduction
Frequency-domain analysis of time series data has been conducted in many applied fields. Central to this analysis is the (power) spectral density, a crucial element whose estimation has garnered significant attention. For instance, in seismology, earthquake source parameters are delineated through the spectral parameter estimation [5, 33]. In geodesy, temporal correlation of noises in the Global Navigation Satellite System is captured by spectral parameter estimation [17]. In audio signal processing, various methodologies have been developed via spectral density estimation to achieve signal separation [20].
Traditional spectral estimation often hinges on the Whittle likelihood maximization [31], which is equivalent to minimizing the Itakura–Saito divergence [13] between a periodogram and a selected class of spectral densities. Broadening the perspective, one can approach such estimations through the lens of spectral divergences, statistical metrics for dissimilarities between two spectral densities. Estimation based on general spectral divergences is comprehensively explored by [28].
In this paper, we take another look at a particular class of spectral divergence: the spectral $\alpha$-Rényi divergences [29, 34, 32, 8, 9]. Notably, the spectral Rényi divergence encompasses the Itakura–Saito divergence as a special instance. In the probabilistic and information-theoretic literature, this class has been mentioned in the extended discussion of “Pinsker’s information theoretic justification of the Itakura-Saito distortion measure” [24]. In the statistical literature, the robustness of the spectral Rényi divergence against spectral peaks is pointed out [34], which has been utilized in clustering analysis of time series [32, 11].
The primary aim of this paper is to elucidate new properties of this class that shed light on the robustness in the spectral parameter estimation. Specifically, we show that (i) the $\gamma$-divergence [7] between probability densities of the processes leads to the spectral Rényi divergence, and (ii) the spectral Rényi divergence has a variational representation. Inspired by these new results, we explore (iii) an additional robustness property of the minimum spectral Rényi divergence estimate. The minimum spectral Rényi divergence estimate has been shown to be robust against time series outliers in the frequency domain [34, 32, 11]. Expanding on these findings, we show that the optimization path in obtaining the minimum spectral Rényi divergence estimate is stable in the presence of time series outliers, a quality that the minimum Itakura–Saito divergence estimator does not have.
Let us explain the implications of our results. The first result bridges spectral and probabilistic divergences, providing an information-theoretic justification of the spectral Rényi divergence. The second result gives a decomposition of spectral Rényi divergence in terms of the Itakura–Saito divergence. The first two results contribute to understanding the robustness of the spectral Rényi divergence. The third result has both theoretical and practical implications. Theoretically, it gives a novel connection between robust spectral analysis (c.f., [19]) and optimization theory. Developing robust statistics from an algorithmic or computational perspective has been a recent topic of discussion (c.f., [6]). Our results provide an additional perspective on robust statistics from a computational viewpoint, focusing on the stability of optimization paths. Practically, outliers in the frequency domain often emerge due to insufficient detrending [10, 12, 22]. Yet, when using the Itakura–Saito divergence, practitioners should exercise greater caution regarding detrending and other pre-processing steps. In comparison, spectral Rényi divergences offer more stable estimation results without necessitating specific meticulous pre-processing.
The structure of this paper is as follows. Section 2 delivers new formulae related to the spectral Rényi divergence (Theorem 1 and Theorem 2). Section 3 discusses the stability of optimization paths of the spectral Rényi divergence minimization (Theorem 3). Section 4 presents thorough numerical studies. Proofs of the results are presented in Appendices.
2 Spectral Rényi divergence
This section introduces a class of spectral Rényi divergences and its properties. Let $\mathcal{S}:=\{S(\cdot):\int_{-\pi}^{\pi}S(\omega)\,\mathrm{d}\omega<\infty,\,S(\omega)>0\,\text{for all $\omega\in(0,\pi)$},S(\omega)=S(-\omega)\}$. Fix $S\in\mathcal{S}$ and $\widetilde{S}\in\mathcal{S}$ arbitrarily. Let $p_{n}(\mathbf{x}_{n})$ and $\widetilde{p}_{n}(\mathbf{x}_{n})$, $\mathbf{x}_{n}\in\mathbb{R}^{n}$, be the $n$-th order probability densities of two zero-mean stationary possibly non-Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively. We assume the sampling rate is equal to 1.
For $\alpha\in(0,1)$, a spectral $\alpha$-Rényi divergence is defined as
$D_{\alpha}[\,S\,:\,\widetilde{S}\,]:=\frac{1}{2\pi(1-\alpha)}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$ (1)
For $\alpha=1$, the spectral $\alpha$-Rényi divergence is defined by using the left-hand limit:
$D_{1}[\,S\,:\,\widetilde{S}\,]$ $:=\lim_{\alpha\to 1^{-}}D_{\alpha}[\,S\,:\,\widetilde{S}\,]$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)^{-1}-1+\log\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)\right\}\,\mathrm{d}\omega,$ (2)
which is equal to the Itakura–Saito divergence [25, 13].
The name of this class comes from the fact that the class is induced by the limit of the probabilistic Rényi divergence between stationary Gaussian processes, which was pointed out by [29]. That is, if two densities $p_{n}$ and $\widetilde{p}_{n}$ are Gaussian, the following holds for $\alpha\in(0,1]$:
$\lim_{n\to\infty}\frac{2}{n}D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=D_{\alpha}[\,S\,:\,\widetilde{S}\,],$
where $D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]$ is the (probabilistic) $\alpha$-Rényi divergence:
$D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]:=\frac{1}{\alpha-1}\log\int_{\mathbb{R}^{n}}\left(\frac{p_{n}(\mathbf{x}_{n})}{\widetilde{p}_{n}(\mathbf{x}_{n})}\right)^{\alpha-1}p_{n}(\mathbf{x}_{n})\,\mathrm{d}\mathbf{x}_{n}.$
The spectral $\alpha$-Rényi divergence with $\alpha\in(0,1]$ is actually a (statistical) divergence: that is, $D_{\alpha}[S,\widetilde{S}]\geq 0$ and $D_{\alpha}[S,\widetilde{S}]=0$ if and only if $S=\widetilde{S}$. This follows since the one-step ahead prediction error variance (the innovation variance) derived from the Kolmogorov–Szëgo formula (Theorem 5.8.1. of [3])
$\sigma^{2}_{S}:=\exp\left(\frac{1}{2\pi}\int_{-\pi}^{\pi}\log{S(\omega)}\,\mathrm{d}\omega\right)$
is log-concave with respect to $S$.
For $\alpha\in(0,1]$, we shall prepare a discrete version of the spectral Rényi divergence as follows:
$D_{\alpha}^{(n)}[\,S\,:\,\widetilde{S}\,]=\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\Big{[}\log\left\{\alpha\widetilde{S}\left(\omega\right)+(1-\alpha)S\left(\omega\right)\right\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}\left(\omega\right)-(1-\alpha)\log S\left(\omega\right)\Big{]},$
where $\Omega_{n}:=\{2\pi(t/n):t=-\lceil n/2\rceil+1,\ldots,-1,0,1,\ldots,\lfloor n/2\rfloor\}$. By the convergence of the Riemann sums, $D^{(n)}_{\alpha}$ converges to $D_{\alpha}$ as $n$ goes to infinity.
2.1 Properties of spectral Rényi divergences
Here we present two new formulae that have not been investigated in the literature: (i) the asymptotic equivalence between the $\gamma$-divergence [7] and the spectral Rényi divergence; and (ii) the variational representation of spectral Rényi divergences.
We first state the asymptotic equivalence between the $\gamma$-divergence and the spectral Rényi divergence. Let $G_{\gamma}[\,p\,:\,q\,]$ be the $\gamma$-divergence between two probability densities $p$ and $q$:
$G_{\gamma}[\,p\,:\,q\,]=\frac{1}{\gamma(1+\gamma)}\log\int(p(x))^{1+\gamma}\,\mathrm{d}x-\frac{1}{\gamma}\log\int p(x)(q(x))^{\gamma}\,\mathrm{d}x$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{1+\gamma}\log\int(q(x))^{1+\gamma}\,\mathrm{d}x.$
Theorem 1 (The $\gamma$-divergence leads to spectral Rényi divergences).
Let $\alpha\in(0,1)$ and $S,\widetilde{S}\in\mathcal{S}$. If two densities $p_{n}$ and $\widetilde{p}_{n}$ are the $n$-th order probability densities of stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, we have
$\lim_{n\to\infty}\frac{2}{n}G_{\gamma}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=D_{\alpha}[\,S\,:\,\widetilde{S}\,],$
where $\gamma=\alpha^{-1}-1>0$.
The proof is given in Appendix A. It employs the direct evaluation of the $\gamma$-divergence and Szegö’s limit theorem[27, 15].
We next state the variational representation of the spectral Rényi divergence.
Theorem 2 (Variational representation of the spectral Rényi divergence).
For $\alpha\in(0,1)$, and $S,\widetilde{S}\in\mathcal{S}$, we have
$D_{\alpha}\,[\,S\,:\,\widetilde{S}\,]$ $=\frac{1}{1-\alpha}\min_{S^{\prime}\in\mathcal{S}}\left\{\alpha\,D_{1}\,[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)\,D_{1}\,[\,S\,:\,S^{\prime}\,]\right\},$ (3)
where the minimum is achieved by $\underline{S}:=\alpha\widetilde{S}+(1-\alpha)S$. The same characterization holds for $D^{(n)}_{\alpha}$.
Proof.
Observe that for any $S^{\prime}\in\mathcal{S}$, we have
$\alpha D_{1}[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)D_{1}[\,S\,:\,S^{\prime}\,]$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\Bigg{\{}\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log S^{\prime}(\omega)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)\Bigg{\}}\,\mathrm{d}\omega$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\Bigg{\{}\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log\frac{S^{\prime}(\omega)}{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}\Bigg{\}}\,\mathrm{d}\omega$
$\,\,+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega$
$=D_{1}[\,\alpha\widetilde{S}+(1-\alpha)S\,:\,S^{\prime}\,]$
$\,\,+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$
So, the minimizer of the left hand side of the above equation is $\alpha\widetilde{S}+(1-\alpha)S$, which proves the assertion. ∎
The dual representation also holds:
Proposition 1.
For $\alpha\in(0,1)$, and $S,\widetilde{S}\in\mathcal{S}$, we have
$D_{\alpha}\,[\,S\,:\,\widetilde{S}\,]$ $=\frac{1}{1-\alpha}\min_{S^{\prime}\in\mathcal{S}}\left\{\alpha\,D_{1}\,[\,S^{\prime}\,:\,S\,]+(1-\alpha)\,D_{1}\,[\,S^{\prime}\,:\,\widetilde{S}\,]\right\},$ (4)
where the minimum is attained by $\overline{S}:=(\alpha S^{-1}+(1-\alpha)\widetilde{S}^{-1})^{-1}$. The same characterization holds for $D^{(n)}_{\alpha}$.
Proof.
Observe that we have, for any $S^{\prime}\in\mathcal{S}$,
$\alpha D_{1}[\,S^{\prime}\,:\,S\,]+(1-\alpha)D_{1}[\,S^{\prime}\,:\,\widetilde{S}\,]$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\left(\frac{\overline{S}(\omega)}{S^{\prime}(\omega)}\right)^{-1}-1+\alpha\log S(\omega)+(1-\alpha)\log\widetilde{S}(\omega)-\log S^{\prime}(\omega)\right]\,\mathrm{d}\omega$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\left(\frac{\overline{S}(\omega)}{S^{\prime}(\omega)}\right)^{-1}-1+\log\frac{\overline{S}(\omega)}{S^{\prime}(\omega)}\right]\,\mathrm{d}\omega$
$\qquad+\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\alpha\log S(\omega)+(1-\alpha)\log\widetilde{S}(\omega)-\log\overline{S}(\omega)\right]\,\mathrm{d}\omega$
$=D_{1}[\,S^{\prime}\,:\,\overline{S}\,]+(1-\alpha)D_{\alpha}\,[\,S\,:\,\widetilde{S}\,].$
So, the minimizer of the left hand side of the above equation is $\overline{S}$, which proves (4). ∎
Remark that this variational representation is a spectral version of the characterization of Rényi divergence in relation to the composite hypothesis testing [26, 30].
2.2 Robustness of the spectral Renyi divergence
A tantalizing feature of the spectral $\alpha$-Rényi divergence for $\alpha\in(0,1)$ is that it is robust with respect to outliers in the frequency domain, which has been discussed in [34, 11]. Theorem 1 explains about this robustness as the $\gamma$-divergence induces the robust parameter estimation scheme [7]. Theorem 2 also provides another explanation about the robustness of the spectral Rényi divergence as discussed in this subsection. Let $\mathcal{S}_{\Theta}:=\{S_{\theta}\in\mathcal{S}:\theta\in\Theta\}$ be a parametric spectral model with a parameter space $\Theta\subset\mathbb{R}^{d}$. Consider the minimum spectral $\alpha$-Rényi divergence estimator $\widehat{\theta}_{\alpha}$ given as
$D^{(n)}_{\alpha}[\,\widetilde{I}_{n}\,:\,S_{\widehat{\theta}_{\alpha}}\,]=\min_{\theta\in\Theta}D^{(n)}_{\alpha}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,],$
where $\widetilde{I}_{n}$ is a nonparametric pilot estimate of the spectral density such as the periodgram
$I_{n}(\omega):=\frac{1}{2\pi n}\left|\sum_{t=1}^{n}x_{t}\mathrm{e}^{-\sqrt{-1}t\omega}\right|^{2}$
and its smoothed versions.
We begin with introducing outliers and contamination in the frequency domain; see Section 8 of [19] for details. Expressing contamination caused by outliers in the frequency domain, we consider a contaminated pilot estimate $\widetilde{I}_{n}^{z,\omega^{*}}$: for $\omega^{*}\in\Omega_{n}$ and $z>0$,
$\widetilde{I}_{n}^{z,\omega^{*}}(\omega)=\begin{cases}\widetilde{I}_{n}(\omega)+z,&\omega=\pm\omega^{*},\\ \widetilde{I}_{n}(\omega),&\omega\neq\pm\omega^{*}.\end{cases}$
This contamination is motivated by the case when periodic components are not suitably subtracted and are contaminated in observed time series [10, 12, 22], say, $x_{t}^{o}=x_{t}+\sqrt{8\pi z/n}\cos(At)$, $t=1,\ldots,n$ with original zero-mean time series $\{x_{t}\}_{t=1,2,\cdots,n}$. In this case, the periodogram of $x_{t}^{o}$ has approximately the form of
$I_{n}^{z,A}(\omega)=I_{n}(\omega)+z(\delta_{A}(\omega)+\delta_{-A}(\omega))+O_{P}(\sqrt{z}),$
where $\delta_{A}(\omega)=1$ if $\omega=A$ and 0 otherwise.
Observe first that the variation of the Itakura–Saito divergence with respect to an outlier in the frequency domain is written as
$D_{1}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-D_{1}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}]=\frac{z}{n}\frac{1}{S_{\theta}(\omega^{*})}-\frac{\log z}{n}+o_{P}(1)$ (5)
with $o_{P}(1)$ being uniform with respect to $z$. For larger value of $z$, this change is influenced by $\theta$, impacting on the behavior of the minimum Itakura–Saito divergence estimator. Now, putting the convex combination of $S_{\theta}$ and $\widetilde{I}^{z,\omega^{*}}_{n}$ into the second slot of the Itakura–Saito divergence mitigates the effect of $z$ as
$D^{(n)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}]-D^{(n)}_{1}[\widetilde{I}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}_{n}]$
$=\frac{1}{n}\frac{\alpha S_{\theta}(\omega^{*})}{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{z}{(1-\alpha)z+\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\quad-\frac{1}{n}\log(1+\widetilde{I}_{n}^{-1}(\omega^{*})z)+\frac{1}{n}\log(1+(1-\alpha)z\{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})\}^{-1})$
$=o_{P}(1)$
with $o_{P}(1)$ being uniform with respect to $z$. This, together with the variational representation (3), implies that
$D_{\alpha}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-D_{\alpha}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}]=\frac{\alpha\log z}{n}+o_{P}(1)$ (6)
with $o_{P}(1)$ being uniform with respect to $z$. So, for large $z$, the change does not vary with respect to $\theta$. This imbues the minimum spectral Rényi divergence estimator with robustness against outliers in the frequency domain.
3 Optimization paths
This section presents a further robustness property of the minimum spectral Rényi divergence estimator. We show that (1) the optimization for the minimum spectral Rényi divergence estimator is robust against to outliers in the frequency domain, whereas (2) the optimization for the minimum Itakura–Saito divergence estimator is sensitive to them.
3.1 Stable optimization path of the Spectral Rényi divergence minimization
We begin with showing that the optimization for the minimum spectral Rényi divergence estimator is robust against to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{\alpha}$: for the $(m+1)$-th step,
$\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]=\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]+\gamma_{m}\,\mathcal{G}_{\alpha}(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n})$
with
$\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ $=-\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}_{n}\,:\,S_{\theta}]$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left[1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega).$ (7)
For theoretical results, we make the following assumptions.
Assumption 1.
We assume the following:
  1. (1)
    The spectral density $S_{\theta}(\omega)$ is differentiable with respect to $\theta$ and we have
    $\sup_{\theta\in\Theta,\,\omega\in[-\pi,\pi]}\max\{\|\nabla_{\theta}S_{\theta}(\omega)\|,\,\|\nabla_{\theta}\log S_{\theta}(\omega)\|,\,\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\}$
    $\qquad\qquad\qquad=:U_{1}<\infty;$
  2. (2)
    The spectral density $S_{\theta}(\omega)$ is twice differentiable with respect to $\theta$ and we have
    $\limsup_{n\to\infty}\left(\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)=:U_{2}<\infty,$
    where let $\|\cdot\|_{\mathrm{op}}$ denote the operator norm.
Assumption 2.
For any $z\geq 0$, $\omega^{*}\in\Omega_{n}$, and $m\in\mathbb{N}$, the $m$-th gradient decent update $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]$ lies in $\Theta$.
Under Assumptions 1 and 2, the subsequent theorem indicates that for any value of $\alpha\in(0,1)$, the path of the gradient descent with a prefixed step size for $\widehat{\theta}_{\alpha}$ is robust against the presence of outliers in the frequency domain.
Theorem 3 (Stability of optimization path with a fixed learning rate sequence).
Fix arbitrary step size sequence $\{\gamma_{k}>0\,:\,k=1,2,\ldots\}$. Under Assumptions 1 and 2, for any $m\in\mathbb{N}$, we have
$\lim_{n\to\infty}\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely}.$ (8)
The proof is given in Appendix B.
Remark 1 (Constant step size).
Assumption 1 ensures that we can obtain an upper bound of the smoothness constant of $D^{(n)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ with respect to $\theta$ uniformly in $z\geq 0$: for any $z\geq 0$, we have
$\sup_{\theta\neq\theta^{\prime}}\frac{\|\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta^{\prime}}]\|}{\|\theta-\theta^{\prime}\|}$
$\leq\sup_{z\geq 0}\sup_{\theta\in\Theta}\|\nabla_{\theta}^{2}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|_{\mathrm{op}}$
$\leq L:=\frac{U_{1}^{2}+U_{2}}{1-\alpha}<\infty,$ (9)
where we use the following expression of the Hessian of $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$:
$\left(-\nabla_{\theta}^{2}D^{(n)}_{\alpha}[\widetilde{I}_{n}:S_{\theta]}]\right)$
$=-\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\frac{\alpha S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\qquad\qquad\frac{(1-\alpha)\widetilde{I}_{n}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\left(\nabla_{\theta}\log S_{\theta}(\omega)\right)\left(\nabla_{\theta}\log S_{\theta}(\omega)\right)^{\top}$
$\quad+\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right\}\nabla^{2}_{\theta}\log S_{\theta}(\omega).$
[…]
$\lim_{m\to\infty}\nabla_{\theta=\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0.$
If the further assumption called the Kurdyka–Łojasiewicz inequality [18, 16] holds at all stationary points, that is, points satisfying $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0$ for the function $\theta\mapsto D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, the bounded gradient descent update converges to a stationary point; c.f., Theorem $3.2$ of [1]. Theorem 3 implies that the finite update sequences for different $z\geq 0$ become the same and all stationary points for different $z\geq 0$ become the same as $n\to\infty$, which gives an new aspect into the robust spectral analysis and the optimization theory.
Remark 2 (AR(1) model).
Here we check Assumption 1 for a first-order autoregressive model. The model spectral density is
$S_{\sigma,\rho}(\omega)=\frac{1}{2\pi}\frac{\sigma^{2}}{1-2\rho\cos(\omega)+\rho^{2}},\quad\sigma>0\quad\text{and}\quad\rho\in(-1,1).$
We reparametrize this by using
$\theta_{1}=\log\sigma\quad\text{and}\quad\theta_{2}=\log\frac{1+\rho}{1-\rho}$
so as to remove constraints in the optimization. Then, the gradient of $\log S_{\theta}(\omega)$ with respect to $\theta$ is given by
$\nabla_{\theta}\log S_{\theta}(\omega)=\begin{pmatrix}2\\ (1+\rho)(1-\rho)\{\cos(\omega)-\rho\}/\{1-2\rho\cos(\omega)+\rho^{2}\}\end{pmatrix}.$
The Hessian of $\log S_{\theta}(\omega)$ is a bit complex but is bounded as
$\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\leq C/(1-|\rho|)^{4}$
for some absolute positive constant $C$. So, Assumption 1 holds for any bounded parameter space $\Theta$ of $\theta$.
Figure 1: Comparison of optimization paths. The optimization paths of the gradient descent for the $AR(1)$ model, with respect to $(\log\sigma,2\{1/(1+\exp(-\rho))-1/2\})$, are presented under the spectral Rényi divergence minimization and the Itakura–Saito divergence minimization. The left figure shows the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $5000$ iterations (the inverted triangles), where the colors indicates different values of the contamination $z$. The right figure shows the corresponding results for the Itakura–Saito divergence minimization.
Theorem 3 establishes a connection between robustness in statistics and optimization theory, offering a novel perspective in robust statistics. While detailed numerical experiments are presented in Section 4, we briefly illustrate the implications through a simple simulation study here.
We consider contaminated observations defined as
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,1000$
with $X_{t}$ following from AR(1) model having $(\sigma,\rho)$. We perform 5000 iterations of gradient descent with a step size of 0.01 with respect to
$(\log\sigma,2\{1/(1+\exp(-\rho))-1/2\}).$
The left panel of Figure 1 illustrates the optimization paths and endpoints based on $D_{\alpha=0.5}[I_{1000}:S_{\theta}]$, starting from the initial value (gray circle). Different colors represent different contamination levels $z\in\{0,10^{2},10^{4}\}$. The figure highlights the robustness of the optimization paths under varying contamination levels. In contrast, the right figure of Figure 1 shows the optimization paths and the stopping points based on the Itakura–Saito divergence $D_{1}[I_{1000}\,:\,S_{\theta}]$ starting from the initial value (the gray circle). This figure demonstrates the non-robustness of the optimization paths of the Itakura–Saito divergence minimization under contamination. The instability observed here will be analyzed further in the subsequent subsection.
Practically, gradient descent updates commonly use line search methods to optimize the step size dynamically. To manage this situation, we present a guiding theorem for gradient descent updates with line search. In this paper, we employ the Armijo condition for this purpose. Consider an objective function $\mathcal{L}(\theta)$ to be minimized and let $\theta^{(m)}$ be the $m$-th gradient descent update for each $m\in\mathbb{N}$. The Armijo condition with $c\in(0,1)$ selects the step size $\gamma$ in the $m$-th update $\theta^{(m)}$ so as to satisfy
$\mathcal{L}\left(\theta^{(m-1)}-\gamma\nabla_{\theta=\theta^{(m-1)}}\mathcal{L}(\theta)\right)\leq\mathcal{L}\left(\theta^{(m-1)}\right)-c\gamma\left\|\nabla_{\theta=\theta^{(m-1)}}\mathcal{L}(\theta)\right\|^{2},$
which ensures the monotone decrease of the objective function along with the gradient descent update. We denote by $\mathcal{A}[c,\theta,\mathcal{L}]$ the Armijo condition with $c\in(0,1)$ for an objective function $\mathcal{L}(\theta)$ in the update from $\theta$.
Proposition 2 (Stability of line search).
Fix $c\in(0,1)$, $\kappa>0$, and $\underline{\gamma}>0$. Fix also the point $\theta\in\Theta$ such that
$\left\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\right\|>\kappa.$
[…]
The proof is given in Appendix C. It employs the descent lemma (c.f., [2]) together with the comparison of gradients.
Remark 3 (Stability of optimization path with the Armijo condition).
Let $\{\gamma_{m}[\widetilde{I}_{n}]:m=1,2,\ldots\}$ and $\{\gamma_{m}[\widetilde{I}^{z,\omega^{*}}_{n}]:m=1,2,\ldots\}$ the step size sequences for $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$ and $D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, respectively. From Proposition 2, we can assume
$\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}|\gamma_{m}[\widetilde{I}_{n}]-\gamma_{m}[\widetilde{I}_{n}^{z,\omega^{*}}]|\to 0\quad\text{as }\quad n\to\infty.$
Under this additional assumption and if we have
$\inf_{k=0,\ldots,m-1}\|\nabla_{\theta=\widehat{\theta}_{\alpha}^{(k)}[\widetilde{I}_{n}]}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\|>0,$
we can have for any $m\in\mathbb{N}$,
$\lim_{n\to\infty}\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely},$
even when we utilize the line search.
3.2 Instable optimization path of the Itakura–Saito divergence minimization
We next show that the optimization for the minimum Itakura–Saito divergence estimator is sensitive to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\gamma_{1},\gamma_{2},\ldots>0$ of $\widehat{\theta}_{1}$: for the $(m+1)$-th step,
$\widehat{\theta}_{1}^{(m+1)}[\widetilde{I}_{n}]=\widehat{\theta}_{1}^{(m)}[\widetilde{I}_{n}]+\gamma_{m}\,\mathcal{G}_{1}(\widehat{\theta}_{1}^{(m)}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n})$
with
$\mathcal{G}_{1}(\theta\,;\,\widetilde{I}_{n})=\frac{1}{n}\sum_{\omega\in\Omega_{n}}\left[\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}-1\right]\nabla_{\theta}\log S_{\theta}(\omega).$
[…]
Proposition 3.
Fix $n\in\mathbb{N}$. Assume that for $\omega^{*}\in\Omega_{n}$, the inequality
$\inf_{\theta\in\Theta}\|(1/S_{\theta}(\omega^{*}))\nabla_{\theta}\log S_{\theta}(\omega^{*})\|>0$
holds and $\theta^{(0)}$ lies in $\Theta$. Then we have
$\sup_{z>0}\left\|\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]\right\|=\infty\,\,\text{almost everywhere}.$ (10)
Further, assume that for any $z>0$ and $\omega^{*}\in\Omega_{n}$, the gradient decent sequence $\{\widehat{\theta}^{(m)}_{1}[\widetilde{I}^{z,\omega^{*}}]\}_{m=1,2\ldots}$ lies in $\Theta$. Then, even when the gradient descent sequence with $\widetilde{I}^{z,\omega^{*}}_{n}$ has a subsequence converging to a stationary point $\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]$ for $D_{1}^{(n)}[\,\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}\,]$ , i.e.
$\nabla_{\theta=\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}D_{1}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}]=0,$
this stationary point $\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]$ is not a stationary point for $D_{1}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]$:
$\sup_{z>0}\left\|\nabla_{\theta=\widehat{\theta}^{(\infty)}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{1}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]\right\|=\infty\,\,\text{almost everywhere}.$
The proof is given in Appendix D.
4 Simulation studies
We present numerical studies employing the Brune spectral model with attenuation that is motivated by seismological studies [33]. We compare five estimation methods:
  • the minimum spectral Rényi divergence estimators with
    $\alpha\in\{0.5,0.75,0.9\}$
    using the smoothed periodogram ($I^{S}_{n}$);
  • the minimum Itakura–Saito divergence estimator employing the periodogram ($I_{n}$); and
  • the minimum Itakura–Saito divergence estimator with the smoothed periodogram ($I_{n}^{S}$).
For each divergence minimization problem, the minimizer is computed via the gradient descent algorithm with a fixed learning rate $\gamma=0.005$ starting from a specified initial point detailed for each model. We stop the optimization when the optimization step reaches 10000 or when the Euclidean norm of the gradient of each divergence evaluated at each updated estimate is less than $10^{-3}$. For smoothing the periodogram, we apply the modified Daniell smoothers twice, where the length of the first time is $3$ and the length of the second time is $5$.
4.1 The Brune spectral model with attenuation
Figure 2: Spectral densities with the estimates plugged-in without any trend. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 3: Spectral densities with the estimates plugged-in with the trigonometric trends. The red dashed lines denote the frequencies at which the outliers are injected. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Table 1: Mean values of biases with standard deviations without any trend. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),1}$0.03 ($\pm$0.05) -0.12 ($\pm$0.07) 0.30 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.17 ($\pm$0.06) 0.54 ($\pm$0.22)
R ($\alpha=0.90$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.20 ($\pm$0.06) 0.90 ($\pm$0.33)
IS with $I_{n}$ $\theta^{(0),1}$ 42.6 ($\pm$2.66) 893 ($\pm$50.5) 48.2 ($\pm$2.99)
IS with $I_{n}^{S}$ $\theta^{(0),1}$ 3372 ($\pm$5318) 67386 ($\pm$106260) 5259 ($\pm$8327)
R ($\alpha=0.50$) $\theta^{(0),2}$0.03 ($\pm$0.05) -0.10 ($\pm$0.07) 0.23 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),2}$ 0.002 ($\pm$0.06) -0.17 ($\pm$0.07) 0.50 ($\pm$0.21)
R ($\alpha=0.90$) $\theta^{(0),2}$ 0.01 ($\pm$0.06) -0.23 ($\pm$0.05) 1.05 ($\pm$0.34)
IS with $I_{n}$ $\theta^{(0),2}$ 0.06 ($\pm$0.10) -0.04 ($\pm$0.13) 0.06 ($\pm$0.17)
IS with $I_{n}^{S}$ $\theta^{(0),2}$0.29 ($\pm$0.30) 1.95 ($\pm$1.89) 1.86 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$0.10 ($\pm$0.04) 0.51 ($\pm$0.13) -0.30 ($\pm$0.06)
R ($\alpha=0.75$) $\theta^{(0),3}$0.05 ($\pm$0.06) 0.08 ($\pm$0.19) 0.05 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),3}$0.007 ($\pm$0.06) -0.15 ($\pm$0.13) 0.72 ($\pm$0.35)
IS with $I_{n}$ $\theta^{(0),3}$ 0.10 ($\pm$0.07) 0.42 ($\pm$0.26) -0.28 ($\pm$0.13)
IS with $I_{n}^{S}$ $\theta^{(0),3}$0.29 ($\pm$0.15) 2.09 ($\pm$1.68) 1.50 ($\pm$0.72)
Table 2: Mean values of biases with standard deviations with the trigonometric trends. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),1}$ 0.22 ($\pm$0.07) -0.22 ($\pm$0.05) 0.27 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),1}$ 0.53 ($\pm$0.08) -0.34 ($\pm$0.04) 0.47 ($\pm$0.19)
R ($\alpha=0.90$) $\theta^{(0),1}$ 1.45 ($\pm$0.10) -0.48 ($\pm$0.03) 0.74 ($\pm$0.31)
IS with $I_{n}$ $\theta^{(0),1}$ 75.8 ($\pm$2.85) 1524 ($\pm$55.86) 60.0 ($\pm$3.04)
IS with $I_{n}^{S}$ $\theta^{(0),1}$ 3405 ($\pm$5318) 68019 ($\pm$106263) 5271 ($\pm$8327)
R ($\alpha=0.50$) $\theta^{(0),2}$ 0.21 ($\pm$0.07) -0.20 ($\pm$0.06) 0.23 ($\pm$0.14)
R ($\alpha=0.75$) $\theta^{(0),2}$ 0.53 ($\pm$0.08) -0.33 ($\pm$0.04) 0.45 ($\pm$0.19)
R ($\alpha=0.90$) $\theta^{(0),2}$ 1.48 ($\pm$0.10) -0.49 ($\pm$0.03) 0.79 ($\pm$0.33)
IS with $I_{n}$ $\theta^{(0),2}$ 3.92 ($\pm$0.07) -0.35 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),2}$ 2.83 ($\pm$0.97) -0.09 ($\pm$0.97) 2.30 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$ 0.08 ($\pm$0.06) 0.24 ($\pm$0.15) -0.25 ($\pm$0.10)
R ($\alpha=0.75$) $\theta^{(0),3}$ 0.46 ($\pm$0.08) -0.27 ($\pm$0.07) 0.26 ($\pm$0.16)
R ($\alpha=0.90$) $\theta^{(0),3}$ 1.40 ($\pm$0.09) -0.47 ($\pm$0.03) 0.69 ($\pm$0.29)
IS with $I_{n}$ $\theta^{(0),3}$ 3.90 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),3}$ 2.80 ($\pm$0.95) -0.13 ($\pm$0.59) 2.01 ($\pm$0.86)
Consider the Brune spectral model [4] with attenuation [14] for velocity waveforms:
$S^{\mathrm{BA}}_{\theta}(\omega):=\omega^{2}\left[\frac{\sigma^{2}}{\{1+(\omega/\omega_{c})^{2}\}^{2}}\right]\exp(-\omega/Q),\;\;\theta=(\sigma,\omega_{c},Q).$
The Brune spectral model $\sigma^{2}/\{1+(\omega/\omega_{c})^{2}\}^{2}$ for displacement waveforms is the spectral density model of Matérn process [21] with the spectral decay rate fixed to 2 and often used in the earthquake source estimation [5, 33]. The attenuation factor $\exp(-\omega/Q)$ represents anelasticity of the medium through which the seismic wave propagates. The multiplicative factor $\omega^{2}$ comes from the time derivative that converts a displacement waveform to a velocity waveform.
For the experiments, we use the true value $\theta^{\ast}:=(\sigma^{\ast},\omega_{c}^{\ast},Q^{\ast})=(1,1,1)$. We generate a time series $\mathbf{x}_{n=2^{10}}$ with its spectral density $S_{\theta^{*}}^{\mathrm{BA}}$ 100 times, and if we consider outliers in the frequency domain, we add two trigonometric trends to the time series as
$x_{t}^{o}=x_{t}+\sqrt{8\pi z_{1}/n}\sin(A_{1}t)+\sqrt{8\pi z_{2}/n}\sin(A_{2}t),\,\,t=1,\ldots,n$
with $z_{1}=z_{2}=100$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$. We report the results based on different values of $z_{1}=z_{2}$ in Appendix E. We use three different initial values: $\theta^{(0),1}=(1,0.1,1)$, $\theta^{(0),2}=(1,1,1)$, and $\theta^{(0),3}=(1,2,1)$. Setting $\sigma^{*}$ and $Q^{*}$ as initial values does not impact on the behavior of the minimum spectral Rényi divergence estimates; we report the results based on different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$ in Appendix E.
Figures 2 and 3 display the estimated spectral densities. In each figure, the following representations are used:
  • The solid gray curve represents the periodogram $I_{n}$;
  • the solid black curve denotes the true spectral density $S^{\mathrm{BA}}_{\theta^{*}}$;
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged-in;
  • the solid red, salmon pink, and green curves display the spectral densities with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged-in, respectively.
For any initial value and regardless of the existence of the trend, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence aligns well with the true spectral density. The minimum spectral Rényi divergence estimate with any $\alpha\in\{0.5,0.75,0.9\}$ remains stable with respect to the choice of the initial value. Conversely, the minimum Itakura–Saito divergence estimate shows sensitivity to this choice.
Tables 1 and 2 summaries the estimation results. Comparing the results based on different initial values, we find that the minimum Itakura–Saito divergence estimates are sensitive to the initial value of the optimization regardless of the existence of the smoother, while the spectral Rényi divergence yields estimation results stable to the choice of the initial value. The minimum Itakura–Saito divergence estimate with the smoothed periodogram performs quite worser than that with the periodogram. In the presence of outliers in the frequency domain, the minimum Rényi divergence estimate with $\alpha=0.5$ performs the best regardless of the choice of the initial value in this example.
5 Discussions
The value of $\alpha$ controls the trade-off between the the efficiency without the presence of outliers and the robustness even for larger trends. For an objective selection of $\alpha$, constructing an information criterion is one possible direction, but requires sufficient discussions and experiments. In practice, monitoring the behaviours of the minimum spectral Rényi divergence estimates for several values of $\alpha$ (say, $\alpha=0.5,0.75,0.9$) would be a workaround.
The minimum spectral Rényi divergence estimates can accommodate a superposition of outliers in the frequency domain, which is omitted in this paper because the extension is straightforward. This is also confirmed by numerical experiments in Section 4.
Numerical experiments presented in Section 4 suggest that the minimum spectral Rényi divergence estimates are robust to the choice of an initial value, while the minimum Itakura–Saito divergence estimate is sensitive to it. These robustness and sensitivity are confirmed in the other examples such as autoregressive models. The structure of the gradient $\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ in the comparison with that of $\mathcal{G}_{1}(\theta\,;\,\widetilde{I}_{n})$ can give a clue to this, and so detailed analysis of these gradient landscapes would be one of interesting research directions.
6 Acknowledgement
The authors would like to thank Akifumi Okuno, Mirai Tanaka, Kei Kobayashi, Yuta Koike, Tomoyuki Higuchi, Hironori Fujisawa, Junho Yang, Masayuki Kano, and Shotaro Yagishita for their comments. This work is supported by JSPS KAKENHI (19K20222, 21H05205, 21K12067, 23K11024), MEXT (JPJ010217), and “Strategic Research Projects” grant (2022-SRP-13) from ROIS (Research Organization of Information and Systems). The simulation code is available at https://github.com/t-tetsuya/Spectral_Renyi.git.
Supplement to “On robustness of Spectral Rényi divergence” by Tetsuya Takabatake and Keisuke Yano
This supplement presents
  • Proof of Theorem 1 in Appendix A;
  • Proof of Theorem 3 in Appendix B;
  • Proof of Proposition 2 in Appendix C;
  • Proof of Proposition 3 in Appendix D; and
  • Additional simulation studies in Appendix E.
Appendix A Proof of Theorem 1
We firstly introduce the notation used in the proof. For each spectral density function $S$ of a stationary time series $\{X_{t}\}_{t\in\mathbb{Z}}$, we write $\Sigma_{n}(S)$ as the $n\times n$-Toeplitz matrix whose $(s,t)$th element is given by
$R(t-s)=\int_{-\pi}^{\pi}e^{\sqrt{-1}(t-s)\omega}S(\omega)\,\mathrm{d}\omega,\ \ s,t=1,\ldots,n,$
which is corresponding to the covariance matrix of $(X_{1},X_{2},\cdots,X_{n})^{\top}$. Then, the probability density function of $(X_{1},\cdots,X_{n})^{\top}$ is expressed by
$p_{S}(\mathbf{x})=\frac{1}{\sqrt{(2\pi)^{n}\det[\Sigma_{n}(S)]}}\exp\left(-\frac{1}{2}\mathbf{x}^{\top}\Sigma_{n}(S)^{-1}\mathbf{x}\right),\ \ \mathbf{x}\in\mathbb{R}^{n}.$ (11)
Let $\gamma\in(0,\infty)$. First introduce the $\gamma$-cross entropy $d_{\gamma}(p_{S},p_{\widetilde{S}})$ between two probability density functions $p_{S}$ and $p_{\widetilde{S}}$, which are not necessarily Gaussian probability density functions, is defined by
$d_{\gamma}[p_{S}:p_{\widetilde{S}}]:=-\frac{1}{\gamma}\log\int_{\mathbb{R}^{n}}p_{S}(\mathbf{x})p_{\widetilde{S}}(\mathbf{x})^{\gamma}\,\mathrm{d}\mathbf{x}+\frac{1}{1+\gamma}\log\int_{\mathbb{R}^{n}}p_{\widetilde{S}}(\mathbf{x})^{1+\gamma}\,\mathrm{d}\mathbf{x}.$
Then the $\gamma$-divergence can be written as $D_{\gamma}[p_{S}:p_{\widetilde{S}}]=d_{\gamma}[p_{S}~{}:p_{\widetilde{S}}]-d_{\gamma}[p_{S}:p_{S}]$.
Now we go back to the proof. We firstly compute the $\gamma$-cross entropy $d_{\gamma}[p_{S}:p_{\widetilde{S}}]$ for Gaussian probability density functions $p_{S}$ and $p_{\widetilde{S}}$, and then the conclusion directly follows from Szegö’s limit theorem [27, 15]. Since $S$ and $\widetilde{S}$ are positive on $(0,\pi)$, we can show that $\Sigma_{n}(S)$ and $\Sigma_{n}(\widetilde{S})$ are positive definite so that $\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)$ is also positive definite for any $\gamma>0$, which yields $\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)$ is invertible. Then we obtain
$\int_{\mathbb{R}^{n}}p_{S}(\mathbf{x})p_{\widetilde{S}}(\mathbf{x})^{\gamma}\,\mathrm{d}\mathbf{x}$ =$=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
$\qquad\quad\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}\mathbf{x}^{\ast}[\Sigma_{n}(S)^{-1}+\gamma\Sigma_{n}(\widetilde{S})^{-1}]\mathbf{x}\right)\,\mathrm{d}\mathbf{x}$
=$=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
$\qquad\quad\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}\mathbf{x}^{\top}\Sigma_{n}(S)^{-1}[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]\Sigma_{n}(\widetilde{S})^{-1}\mathbf{x}\right)\,\mathrm{d}\mathbf{x}$
=$=\left\{\frac{\pi^{n}\det[\Sigma_{n}(\widetilde{S})(\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S))^{-1}\Sigma_{n}(S)]}{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}}\right\}^{\frac{1}{2}}$
=$=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]^{-1}\right\}^{\frac{1}{2}}.$
Similarly, we also obtain
$\int_{\mathbb{R}^{n}}p_{\widetilde{S}}(\mathbf{x})^{1+\gamma}\,\mathrm{d}\mathbf{x}$ =$=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[(1+\gamma)\Sigma_{n}(\widetilde{S})]^{-1}\right\}^{\frac{1}{2}}$
=$=\left\{\pi^{-n\gamma}(1+\gamma)^{-n}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right\}^{\frac{1}{2}}.$
Combining these yields
$2d_{\gamma}[p_{S}:p_{\widetilde{S}}]=$ $-\frac{1}{\gamma}\log\left[\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]^{-1}\right]$
$\qquad+\frac{1}{1+\gamma}\log\left[\pi^{-n\gamma}(1+\gamma)^{-p}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right]$
$=$ $n\log(\pi)-\frac{n}{1+\gamma}\log[(1+\gamma)\pi^{\gamma}]$
$\qquad+\left[\frac{1}{\gamma}\log\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]-\frac{1}{\gamma(1+\gamma)}\log\det[\Sigma_{n}(\widetilde{S})]\right].$
Putting $S$ into $\widetilde{S}$, we obtain
$2d_{\gamma}[p_{S}:p_{S}]$ $=n\log(\pi)-\frac{n}{1+\gamma}\log[(1+\gamma)\pi^{\gamma}]+\frac{1}{\gamma}\log\det[(1+\gamma)\Sigma_{n}(S)]$
$\qquad\qquad\qquad\qquad\qquad-\frac{1}{\gamma(1+\gamma)}\log\det[\Sigma_{n}(S)]$
$=n\log(\pi)-\frac{n}{1+\gamma}\log[(1+\gamma)\pi^{\gamma}]+\frac{n}{\gamma}\log(1+\gamma)+\frac{1}{1+\gamma}\log\det[\Sigma_{n}(S)].$
Thus we get
$2D_{\gamma}[p_{S}:p_{\widetilde{S}}]$ $=\frac{1}{\gamma}\Bigg{[}\log\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]-\frac{1}{1+\gamma}\log\det[\Sigma_{n}(\widetilde{S})]$
$\qquad\qquad\qquad\qquad\qquad\qquad-\log(1+\gamma)^{p}-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]\Bigg{]}$
$=\frac{1}{\gamma}\Bigg{[}\log\det\left[\frac{\gamma}{1+\gamma}\Sigma_{n}(S)+\frac{1}{1+\gamma}\Sigma_{n}(\widetilde{S})\right]$
$\qquad\qquad\qquad\qquad-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]-\frac{1}{1+\gamma}\log\det[\Sigma_{n}(\widetilde{S})]\Bigg{]}$
$=\frac{1}{\gamma}\left[\log\det[\Sigma_{n}(\bar{S}^{(\gamma)})]-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]-\frac{1}{1+\gamma}\log\det[\Sigma_{n}(\widetilde{S})]\right],$
where $\bar{S}^{(\gamma)}(\omega):=\frac{\gamma}{1+\gamma}S(\omega)+\frac{1}{1+\gamma}\widetilde{S}(\omega)$. Using Szegö’s limit theorem [27, 15], we conclude
$\lim_{n\to\infty}\frac{2}{n}D_{\gamma}[p_{S}:p_{\widetilde{S}}]$ $=\frac{1}{2\pi\gamma}\int_{-\pi}^{\pi}\left[\log{\bar{S}^{(\gamma)}(\omega)}-\frac{\gamma}{1+\gamma}\log{S(\omega)}-\frac{1}{1+\gamma}\log{\widetilde{S}(\omega)}\right]\,\mathrm{d}\omega$
$=D_{\gamma}[S:\widetilde{S}].$
This completes the proof.
Appendix B Proof of Theorem 3
Proof.
We prove Theorem 3 by the mathematical induction.
Step 1: Consider $m=1$. Observe
$\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]$
=$=\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\right\}-\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]-\theta^{(0)}\right\}$
$=\frac{\gamma_{1}\alpha}{(1-\alpha)n}\Bigg{(}\frac{S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\qquad\qquad\qquad\qquad-\frac{S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)(\widetilde{I}_{n}(\omega^{*})+z)}\Bigg{)}\nabla_{\theta^{(0)}}\log S_{\theta}(\omega^{*})$
$=\frac{\gamma_{1}}{(1-\alpha)n}\frac{\alpha S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{(1-\alpha)z\nabla_{\theta^{(0)}}S_{\theta}(\omega^{*})}{(1-\alpha)z+(1-\alpha)\widetilde{I}_{n}(\omega^{*})+\alpha S_{\theta^{(0)}}(\omega^{*})}.$
This yields
$\left\|\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]\right\|\leq\frac{1}{n}\frac{\gamma_{1}}{1-\alpha}\|\nabla_{\theta^{(0)}}S_{\theta}(\omega^{*})\|,$ (12)
which proves the assertion for $m=1$.
Step 2: Let $m\in\mathbb{N}>1$. Assume that for any $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for $n\geq N$, we have
$\|\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\|\leq\varepsilon.$
By the cancelling technique and by the triangle inequality, we get
$\left\|\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]\right\|$ =$=\Bigg{\|}\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]\right\}$
$\qquad-\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\right\}-\left\{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\right\}\Bigg{\|}$
$\leq\gamma_{m}\left\|\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)\right\|+\varepsilon.$
Observe that by applying the Taylor theorem to $\nabla_{\theta}\log S_{\theta}(\omega)$ and letting $t$ be some point on some line connecting $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]$ to $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]$, we get
$\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega)$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{t}^{2}\log S_{\theta}(\omega)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}])$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
where $O(1)$ in the rightmost side implies a vector whose norm is $O(1)$. This yields
$\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\Bigg{\{}\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\qquad\qquad-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\Bigg{\}}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)$
$=\frac{\alpha}{(1-\alpha)n}(T_{1}+T_{2})+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
where
$T_{1}$ $:=\sum_{\omega\in\Omega_{n}}\Bigg{\{}\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\quad-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\Bigg{\}}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega),$
$T_{2}$ $:=\sum_{\omega\in\Omega_{n}}\Bigg{\{}\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}$
$\qquad\qquad\quad-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\Bigg{\}}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega).$
Here we have
$T_{1}$ $=\frac{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{(1-\alpha)z}{(1-\alpha)z+\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\qquad\frac{\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega^{*})}{\alpha}$
$=O\left(\frac{1}{\alpha}\sup_{\theta,\,\omega}\|\nabla_{\theta}\log S_{\theta}(\omega)\|\right)$
and
$T_{2}$ $=\sum_{\omega\in\Omega_{n}}\frac{(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}$
$\qquad\qquad\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)-S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$=O\left(\frac{n\varepsilon}{\alpha}\sup_{\theta,\theta^{\prime},\omega}\|\nabla_{\theta^{\prime}}S_{\theta}(\omega)\|\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\right).$
Then we obtain
$\left\|\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)\right\|$
$\leq\frac{1}{(1-\alpha)n}\sup_{\theta,\omega}\|\nabla_{\theta}\log S_{\theta}(\omega)\|$
$\,\,+\varepsilon\Bigg{\{}\frac{1}{(1-\alpha)}\Big{(}\sup_{\theta,\omega}\|\nabla_{\theta,\omega}S_{\theta}(\omega)\|\Big{)}^{2}+\frac{2}{(1-\alpha)}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\Bigg{\}}$
and thus
$\left\|\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]\right\|$
$\leq\frac{\gamma_{m}}{(1-\alpha)n}\sup_{\theta,\omega}\|\nabla_{\theta}\log S_{\theta}(\omega)\|$
$\qquad+\varepsilon\Bigg{\{}1+\frac{\gamma_{m}}{(1-\alpha)}\Big{(}\sup_{\theta,\omega}\|\nabla_{\theta,\omega}S_{\theta}(\omega)\|\Big{)}^{2}+\frac{2\gamma_{m}}{(1-\alpha)}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\Bigg{\}}.$
This implies that the assertion holds for $m+1$ and by the mathematical induction, we get the conclusion. ∎
Appendix C Proof of Proposition 2
Proof.
Let $\gamma$ be the step size satisfying the Armijo condition $\mathcal{A}[c\,,\,\theta\,,\,D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]]$ and $\gamma\leq 2(1-c)/L$. Let $\mathcal{P}_{\gamma}[\theta]$ be
$\mathcal{P}_{\gamma}[\theta]:=\theta-\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}].$
We begin with utilizing the descent lemma (c.f., [2]):
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]$ $\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]+\langle\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\,,\,\mathcal{P}_{\gamma}[\theta]-\theta\rangle$
$\quad+\frac{L}{2}\|\mathcal{P}_{\gamma}[\theta]-\theta\|^{2},$ (13)
[…]
$\left|\left\langle v\,,\,(\mathcal{P}_{\gamma}[\theta]-\theta)\right\rangle-\left\langle v\,,\,\left(-\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\right)\right\rangle\right|$ $\leq\frac{1}{n}\frac{\gamma}{1-\alpha}\|v\|\|\nabla_{\theta}S_{\theta}(\omega^{*})\|$
$\leq\frac{1}{n}\frac{2U_{1}}{L(1-\alpha)}\|v\|,$
where $U_{1}$ is defined in Assumption 1 and $L$ is defined in (9). Putting the quantity $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ into $v$ in the above inequality, we get
$\langle\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\,,\,\mathcal{P}_{\gamma}[\theta]-\theta\rangle\leq-\gamma\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\frac{1}{n}\frac{2U_{1}^{2}}{L(1-\alpha)^{2}},$ (14)
where we use the inequality
$\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|\leq\frac{\alpha U_{1}}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left|\frac{(\alpha-1)S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}^{z,\omega^{*}}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}^{z,\omega^{*}}(\omega)}\right|\leq\frac{U_{1}}{1-\alpha};$
see (7) for the expression of $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$. Equation (12) also gives
$\left\|\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]-\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\right\|\leq\frac{\gamma U_{1}}{n(1-\alpha)}\leq\frac{2U_{1}}{nL(1-\alpha)}$
so that we obtain
$\|\mathcal{P}_{\gamma}[\theta]-\theta\|^{2}\leq\gamma^{2}\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\left(\frac{2}{n}+\frac{1}{n^{2}}\right)\frac{4U_{1}^{2}}{L^{2}(1-\alpha)^{2}}.$ (15)
Together with (13), equations (14) and (15) imply
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]-\left(\gamma-\frac{L}{2}\gamma^{2}\right)\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\varepsilon_{n},$
where $\varepsilon_{n}$ is given by
$\varepsilon_{n}:=\left(\frac{3}{n}+\frac{1}{n^{2}}\right)\frac{2U_{1}^{2}}{L(1-\alpha)^{2}}.$
For $\gamma$ satisfying $1-(L/2)\gamma\geq c$, we have
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]$ $\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]-c\gamma\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\varepsilon_{n}.$
Together with (15), the assumption $\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\|\geq\kappa$ implies that
$\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}\geq\kappa-\varepsilon_{n}/\underline{\gamma}^{2}.$
Thus, for sufficiently large $n$, we have $\kappa-\varepsilon_{n}/\underline{\gamma}^{2}>0$ so that
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]$
$\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]-\left(c-\frac{\varepsilon_{n}}{\underline{\gamma}(\kappa-\varepsilon_{n}/\underline{\gamma}^{2})}\right)\gamma\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2},$
which concludes the proof. ∎
Appendix D Proof of Proposition 3
Proof.
The difference between $\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]$ and $\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]$ is explicitly written as
$\|\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]\|$ =$=\|\{\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\}-\{\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]-\theta^{(0)}\}\|$
$=\frac{\gamma_{1}z}{n}\left\|\frac{\nabla_{\theta^{(0)}}\log S_{\theta}(\omega^{*})}{S_{\theta^{(0)}}(\omega^{*})}\right\|,$
which proves the first assertion. For the second assertion, observe
$\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg{|}_{\theta=\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}$
$=\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg{|}_{\theta=\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}-z\frac{\nabla_{\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega^{*})}{S_{\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega^{*})}.$
This gives the second assertion, which completes the proof. ∎
Appendix E Additional simulation studies
This appendix presents additional numerical experiments.
Tables 3 and 4 show the estimation results for the Brune spectral model with attenuation in Section 4.1 on the basis of the different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$. For almost all the settings of initial values, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence performs the best.
Table 3: The mean values of biases with standard deviations without any trend. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init. Values greater than 10 are denoted by *.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R($\alpha=0.50$) $\theta^{(0),4}$0.03 ($\pm$0.05) -0.11 ($\pm$0.07) 0.26 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),4}$ 0.04 ($\pm$0.06) -0.30 ($\pm$0.03) 1.33 ($\pm$0.17)
R ($\alpha=0.90$) $\theta^{(0),4}$ 0.09 ($\pm$0.08) -0.39 ($\pm$0.03) 6.13 ($\pm$0.04)
IS with $I_{n}$ $\theta^{(0),4}$
IS with $I_{n}^{S}$ $\theta^{(0),4}$
R ($\alpha=0.50$) $\theta^{(0),5}$0.01($\pm$0.06) -0.18 ($\pm$0.05) 0.47 ($\pm$0.16)
R ($\alpha=0.75$) $\theta^{(0),5}$ 0.01 ($\pm$ 0.06) -0.21 ($\pm$0.05) 0.68 ($\pm$0.22)
R ($\alpha=0.90$) $\theta^{(0),5}$ 0.02 ($\pm$0.06) -0.24 ($\pm$0.04) 1.18 ($\pm$0.34)
IS with $I_{n}$ $\theta^{(0),5}$ 0.07 ($\pm$0.11) -0.08 ($\pm$0.12) 0.12 ($\pm$0.18)
IS with $I_{n}^{S}$ $\theta^{(0),5}$0.28 ($\pm$0.16) 1.84 ($\pm$1.66) 1.90 ($\pm$0.63)
Table 4: The mean values of biases with standard deviations with the trigonometric trends. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),4}$ 0.21 ($\pm$0.07) -0.21 ($\pm$0.06) 0.24 ($\pm$0.14)
R ($\alpha=0.75$) $\theta^{(0),4}$ 0.61 ($\pm$0.09) -0.43 ($\pm$0.02) 1.14 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),4}$ 1.65 ($\pm$0.12) -0.60 ($\pm$0.02) 6.05 ($\pm$0.05)
IS with $I_{n}$ $\theta^{(0),4}$
IS with $I_{n}^{S}$ $\theta^{(0),4}$
R ($\alpha=0.50$) $\theta^{(0),5}$ 0.23 ($\pm$0.07) -0.26 ($\pm$0.04) 0.41 ($\pm$0.16)
R ($\alpha=0.75$) $\theta^{(0),5}$ 0.55 ($\pm$0.08) -0.35 ($\pm$0.04) 0.55 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),5}$ 1.50 ($\pm$0.10) -0.50 ($\pm$0.03) 0.85 ($\pm$0.35)
IS with $I_{n}$ $\theta^{(0),5}$ 3.91 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),5}$ 2.85 ($\pm$0.95) -0.17 ($\pm$0.51) 2.26 ($\pm$0.93)
Figures 4 and 5 display an instance of a set of the estimated spectral densities. In each figure, the following representations are used:
  • The solid gray curve represents the periodogram $I_{n}$;
  • the solid black curve denotes the true spectral density $S^{\mathrm{BA}}_{\theta^{*}}$;
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged-in;
  • the solid red, salmon pink, and green curves display the spectral density with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged-in, respectively.
For any initial value and for any strength of trends, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence performs the best.
Figure 4: Spectral densities with the estimates plugged-in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=2.5$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 5: Spectral densities with the estimates plugged-in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=25$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.