### v4

# On robustness of spectral Rényi divergence

Abstract: This paper studies a specific class of statistical divergences for spectral densities of time series: the spectral $\alpha$-Rényi divergences, which include the Itakura–Saito divergence as a limiting case. The aim of this paper is to highlight both information-theoretic and statistical properties of spectral $\alpha$-Rényi divergences. We reveal the connection between the spectral $\alpha$-Rényi divergence and the $\gamma$-divergence in robust statistics, and a variational representation of the spectral $\alpha$-Rényi divergence. Inspired by these results suggesting “robustness” of spectral $\alpha$-Rényi divergence, we show that the minimum spectral Rényi divergence estimate has a stable optimization path with respect to outliers in the frequency domain, unlike the minimum Itakura–Saito divergence estimator, and thus it delivers more stable estimates, reducing the need for intricate pre-processing.
Keywords: Frequency domain analysis; Optimization theory; Robust statistics; Spectral density; Statistical divergence; Time series analysis.
1 Introduction
Frequency-domain analysis of time series data has been conducted in many applied fields. Central to this analysis is the (power) spectral density, a crucial element whose estimation has garnered significant attention. For instance, in seismology, earthquake source parameters are delineated through the spectral parameter estimation Calderoni and Abercrombie (2023); Yoshimitsu et al. (2023). In geodesy, temporal correlation of noises in the Global Navigation Satellite System is captured by spectral parameter estimation Langbein (2004). In audio signal processing, various methodologies have been developed via spectral density estimation to achieve signal separation Martin (2005).
Traditional spectral estimation often hinges on the Whittle likelihood maximization Whittle (1953), which is equivalent to minimizing the Itakura–Saito divergence Itakura and Saito (1968) between a periodogram and a selected class of spectral densities. Broadening the perspective, one can approach such estimations through the lens of spectral divergences, statistical metrics for dissimilarities between two spectral densities. Estimation based on general spectral divergences is comprehensively explored by Taniguchi (1987).
In this paper, we take another look at a particular class of spectral divergences: the spectral $\alpha$-Rényi divergences Vajda (1989); Zhang and Taniguchi (1995); Kakizawa et al. (1998); Gil et al. (2013); Grivel et al. (2021). Notably, a limit of the spectral Rényi divergence induces the Itakura–Saito divergence. In the probabilistic and information-theoretic literature, this class has been mentioned in the extended discussion of “Pinsker’s information-theoretic justification of the Itakura-Saito distortion measure” Parzen (1993). In the statistical literature, the robustness of the spectral Rényi divergence against spectral peaks is pointed out Zhang and Taniguchi (1995), which has been utilized in clustering analysis of time series Kakizawa et al. (1998); Hirukawa (2006).
The primary aim of this paper is to elucidate new properties of this class that shed light on the robustness in the spectral parameter estimation. Specifically, we show that (i) the $\gamma$-divergence Fujisawa and Eguchi (2008) between probability densities of the processes leads to the spectral Rényi divergence, and (ii) the spectral Rényi divergence has a variational representation. Inspired by these new results, we explore (iii) an additional robustness property of the minimum spectral Rényi divergence estimate. The minimum spectral Rényi divergence estimate has been shown to be robust against time series outliers in the frequency domain Zhang and Taniguchi (1995); Kakizawa et al. (1998); Hirukawa (2006). Expanding on these findings, we show that the optimization path in obtaining the minimum spectral Rényi divergence estimate is stable in the presence of time series outliers, a quality that the minimum Itakura–Saito divergence estimator does not have.
Let us explain the implications of our results. The first result bridges spectral and probabilistic divergences, providing an information-theoretic justification of the spectral Rényi divergence. The second result gives a decomposition of spectral Rényi divergence in terms of the Itakura–Saito divergence. The first two results contribute to understanding the robustness of the spectral Rényi divergence. The third result has both theoretical and practical implications. Theoretically, it gives a novel connection between robust spectral analysis (c.f., Maronna et al. (2019)) and optimization theory. Developing robust statistics from an algorithmic or computational perspective has been a recent topic of discussion (c.f., Diakonikolas and Kane (2023)). Our results provide an additional perspective on robust statistics from a computational viewpoint, focusing on the stability of optimization paths. Practically, outliers in the frequency domain often emerge due to insufficient detrending Heyde and Dai (1996); Iacone (2010); McCloskey and Perron (2013). Yet, when using the Itakura–Saito divergence, practitioners should exercise greater caution regarding detrending and other pre-processing steps. In comparison, spectral Rényi divergences offer more stable estimation results without necessitating elaborate pre-processing. This aspect is further supported by an applied study Kano et al. (2025).
The structure of this paper is as follows. Section 2 delivers new formulae related to the spectral Rényi divergence (Theorems 1 and 2). Section 3 discusses the stability of optimization paths of the spectral Rényi divergence minimization (Theorem 3). Section 4 presents thorough numerical studies. Proofs of the results are presented in Appendices.
2 Spectral Rényi divergence
This section introduces a class of spectral Rényi divergences and their properties. Let $\mathcal{S}:=\{S:[-\pi,\pi]\to[0,\infty]\mid\int_{-\pi}^{\pi}S(\omega)\,\mathrm{d}\omega<\infty,\,S(\omega)>0\,\mbox{and}\,S(\omega)=S(-\omega)\,\text{for all $\omega\in(0,\pi)$}\}$. Fix $S\in\mathcal{S}$ and $\widetilde{S}\in\mathcal{S}$ arbitrarily. Let $p_{n}$ and $\widetilde{p}_{n}$ denote the joint densities of $n$ consecutive observations from two zero-mean stationary processes, possibly non-Gaussian, with spectral densities $S$ and $\widetilde{S}$, respectively. Without loss of generality, we normalize the time scale so that observations are taken at unit time intervals.
For $\alpha\in(0,1)$, a spectral $\alpha$-Rényi divergence is defined as
$D_{\alpha}[\,S\,:\,\widetilde{S}\,]:=\frac{1}{2\pi(1-\alpha)}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$ (1)
For $\alpha=1$, we define $D_{1}[\,S\,:\,\widetilde{S}\,$] by continuity as the limit $\alpha\to 1^{-}$. In this limit, the spectral $\alpha$-Rényi divergence converges to the Itakura–Saito divergence Pinsker (1964); Itakura and Saito (1968):
$\lim_{\alpha\to 1^{-}}D_{\alpha}[\,S\,:\,\widetilde{S}\,$] $=D_{\mathrm{IS}}[\,S\,:\,\widetilde{S}\,$]
$:=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left\{\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)^{-1}-1+\log\left(\frac{\widetilde{S}(\omega)}{S(\omega)}\right)\right\}\,\mathrm{d}\omega.$ (2)
The name of this class stems from the fact that the class is induced by the limit of the probabilistic Rényi divergence between stationary Gaussian processes, as pointed out by Vajda (1989). More precisely, if $p_{n}$ and $\widetilde{p}_{n}$ are the joint densities of $n$ consecutive observations of two zero-mean stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively, then for $\alpha\in(0,1]$, we have
$\lim_{n\to\infty}\frac{2}{n}D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=D_{\alpha}[\,S\,:\,\widetilde{S}\,$],
where $D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]$ is the (probabilistic) $\alpha$-Rényi divergence:
$D_{\alpha}[\,p_{n}\,:\,\widetilde{p}_{n}\,]:=\frac{1}{\alpha-1}\log\int_{\mathbb{R}^{n}}\left(\frac{p_{n}(\mathbf{x}_{n})}{\widetilde{p}_{n}(\mathbf{x}_{n})}\right)^{\alpha-1}p_{n}(\mathbf{x}_{n})\,\mathrm{d}\mathbf{x}_{n}.$
For $\alpha=1$, the probabilistic Rényi divergence is defined by continuity as the limit $\alpha\to 1^{-}$ and coincides with the Kullback–Leibler divergence:
$D_{1}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=\int_{\mathbb{R}^{n}}p_{n}(\mathbf{x}_{n})\log\frac{p_{n}(\mathbf{x}_{n})}{\widetilde{p}_{n}(\mathbf{x}_{n})}\,\mathrm{d}\mathbf{x}_{n}.$
The spectral $\alpha$-Rényi divergence is actually a (statistical) divergence for any $\alpha\in(0,1)$: that is, $D_{\alpha}[\,S\,:\,\widetilde{S}\,]\geq 0$ and $D_{\alpha}[\,S\,:\,\widetilde{S}\,]=0$ if and only if $S=\widetilde{S}$. This follows since the one-step ahead prediction error variance (the innovation variance) derived from the Kolmogorov–Szegö formula (Theorem 5.8.1. of Brockwell and Davis (1991))
$\sigma^{2}_{S}:=\exp\left(\frac{1}{2\pi}\int_{-\pi}^{\pi}\log{S(\omega)}\,\mathrm{d}\omega\right)$
is log-concave with respect to $S$.
For $\alpha\in(0,1)$, we shall prepare a discrete version of the spectral Rényi divergence as follows:
$D_{\alpha}^{(n)}[\,S\,:\,\widetilde{S}\,]=\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\Big[\log\left\{\alpha\widetilde{S}\left(\omega\right)+(1-\alpha)S\left(\omega\right)\right\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}\left(\omega\right)-(1-\alpha)\log S\left(\omega\right)\Big],$
where $\Omega_{n}:=\{2\pi(t/n):t=-\lceil n/2\rceil+1,\ldots,-1,0,1,\ldots,\lfloor n/2\rfloor\}$. By the convergence of the Riemann sums, $D^{(n)}_{\alpha}$ converges to $D_{\alpha}$ as $n$ goes to infinity. The discrete version $D^{(n)}_{\mathrm{IS}}$ of the Itakura–Saito divergence is defined in the same way.
2.1 Properties of spectral Rényi divergences
Here we present two new formulae that have not been investigated in the literature: (i) the asymptotic equivalence between the time-domain $\gamma$-divergence Fujisawa and Eguchi (2008) and the spectral Rényi divergence; and (ii) the variational representation of spectral Rényi divergences.
We first state the asymptotic equivalence between the time-domain $\gamma$-divergence and the spectral Rényi divergence. Let $G_{\gamma}[\,p\,:\,q\,]$ be the $\gamma$-divergence between two probability densities $p$ and $q$:
$G_{\gamma}[\,p\,:\,q\,]=\frac{1}{\gamma(1+\gamma)}\log\int(p(x))^{1+\gamma}\,\mathrm{d}x-\frac{1}{\gamma}\log\int p(x)(q(x))^{\gamma}\,\mathrm{d}x$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{1+\gamma}\log\int(q(x))^{1+\gamma}\,\mathrm{d}x.$
Theorem 1 (The time-domain $\gamma$-divergence leads to the spectral Rényi divergence).
Let $\alpha\in(0,1)$ and $S,\widetilde{S}\in\mathcal{S}$. Let $p_{n}$ and $\widetilde{p}_{n}$ denote the joint densities of $n$ consecutive observations from zero-mean stationary Gaussian processes with spectral densities $S$ and $\widetilde{S}$, respectively. Then, we have
$\lim_{n\to\infty}\frac{2}{n}G_{\gamma}[\,p_{n}\,:\,\widetilde{p}_{n}\,]=D_{\alpha}[\,S\,:\,\widetilde{S}\,$],
where $\gamma=\alpha^{-1}-1>0$.
The proof is given in Appendix A. It employs the direct evaluation of the $\gamma$-divergence and Szegö’s limit theorem Szegö (1920); Kolmogorov (1941).
We next state the variational representation of the spectral Rényi divergence.
Theorem 2 (Variational representation of the spectral Rényi divergence).
For $\alpha\in(0,1)$, and $S,\widetilde{S}\in\mathcal{S}$, we have
$D_{\alpha}\,[\,S\,:\,\widetilde{S}\,$] $=\frac{1}{1-\alpha}\min_{S^{\prime}\in\mathcal{S}}\left\{\alpha\,D_{\mathrm{IS}}\,[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)\,D_{\mathrm{IS}}\,[\,S\,:\,S^{\prime}\,]\right\},$ (3)
where the minimum is achieved by $\underline{S}:=\alpha\widetilde{S}+(1-\alpha)S$. The same characterization holds for $D^{(n)}_{\alpha}$.
Proof.
Observe that for any $S^{\prime}\in\mathcal{S}$, we have
$\alpha D_{\mathrm{IS}}[\,\widetilde{S}\,:\,S^{\prime}\,]+(1-\alpha)D_{\mathrm{IS}}[\,S\,:\,S^{\prime}\,]$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\Bigg\{\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log S^{\prime}(\omega)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)\Bigg\}\,\mathrm{d}\omega$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\Bigg\{\frac{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}{S^{\prime}(\omega)}-1+\log\frac{S^{\prime}(\omega)}{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)}\Bigg\}\,\mathrm{d}\omega$
$\,\,+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega$
$=D_{\mathrm{IS}}[\,\alpha\widetilde{S}+(1-\alpha)S\,:\,S^{\prime}\,]$
$\,\,+\frac{1}{2\pi}\int_{-\pi}^{\pi}[\log\{\alpha\widetilde{S}(\omega)+(1-\alpha)S(\omega)\}-\alpha\log\widetilde{S}(\omega)-(1-\alpha)\log S(\omega)]\,\mathrm{d}\omega.$
So, the minimizer of the left hand side of the above equation is $\alpha\widetilde{S}+(1-\alpha)S$, which proves the assertion. ∎
The dual representation also holds:
Proposition 1.
For $\alpha\in(0,1)$, and $S,\widetilde{S}\in\mathcal{S}$, we have
[…]
where the minimum is attained by $\overline{S}:=(\alpha S^{-1}+(1-\alpha)\widetilde{S}^{-1})^{-1}$. The same characterization holds for $D^{(n)}_{\alpha}$.
Proof.
Observe that we have, for any $S^{\prime}\in\mathcal{S}$,
$\alpha D_{\mathrm{IS}}[\,S^{\prime}\,:\,S\,]+(1-\alpha)D_{\mathrm{IS}}[\,S^{\prime}\,:\,\widetilde{S}\,$]
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\left(\frac{\overline{S}(\omega)}{S^{\prime}(\omega)}\right)^{-1}-1+\alpha\log S(\omega)+(1-\alpha)\log\widetilde{S}(\omega)-\log S^{\prime}(\omega)\right]\,\mathrm{d}\omega$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\left(\frac{\overline{S}(\omega)}{S^{\prime}(\omega)}\right)^{-1}-1+\log\frac{\overline{S}(\omega)}{S^{\prime}(\omega)}\right]\,\mathrm{d}\omega$
$\qquad+\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\alpha\log S(\omega)+(1-\alpha)\log\widetilde{S}(\omega)-\log\overline{S}(\omega)\right]\,\mathrm{d}\omega$
$=D_{\mathrm{IS}}[\,S^{\prime}\,:\,\overline{S}\,]+(1-\alpha)D_{\alpha}\,[\,S\,:\,\widetilde{S}\,$].
So, the minimizer of the left hand side of the above equation is $\overline{S}$, which proves (4). ∎
Note that this variational representation is a spectral version of the characterization of Rényi divergence in relation to the composite hypothesis testing Shayevitz (2010); van Erven and Harremoës (2014).
2.2 Robustness of the spectral Rényi divergence
A tantalizing feature of the spectral $\alpha$-Rényi divergence is that it is robust with respect to outliers in the frequency domain, which has been discussed in Zhang and Taniguchi (1995); Hirukawa (2006). Theorem 1 explains this robustness as the time-domain $\gamma$-divergence induces the robust parameter estimation scheme Fujisawa and Eguchi (2008). Theorem 2 also provides another explanation about the robustness of the spectral Rényi divergence as discussed in this subsection. Let $\mathcal{S}_{\Theta}:=\{S_{\theta}\in\mathcal{S}:\theta\in\Theta\}$ be a parametric spectral model with a parameter space $\Theta\subset\mathbb{R}^{d}$, where $\Theta$ is compact with nonempty interior.
Consider the minimum spectral $\alpha$-Rényi divergence estimator $\widehat{\theta}_{\alpha}$ given as
$D^{(n)}_{\alpha}[\,\widetilde{I}_{n}\,:\,S_{\widehat{\theta}_{\alpha}}\,]=\min_{\theta\in\Theta}D^{(n)}_{\alpha}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,],$
where $\widetilde{I}_{n}$ is a nonparametric pilot estimate of the spectral density such as the periodogram
$I_{n}(\omega):=\frac{1}{2\pi n}\left|\sum_{t=1}^{n}x_{t}\mathrm{e}^{-\sqrt{-1}t\omega}\right|^{2}$
and its smoothed versions.
We begin with introducing outliers and contamination in the frequency domain; see Section 8 of Maronna et al. (2019) for details. Expressing contamination caused by outliers in the frequency domain, we consider a contaminated pilot estimate $\widetilde{I}_{n}^{z,\omega^{*}}$: for $\omega^{*}\in\Omega_{n}$ and $z>0$,
$\widetilde{I}_{n}^{z,\omega^{*}}(\omega)=\begin{cases}\widetilde{I}_{n}(\omega)+z,&\omega=\pm\omega^{*},\\ \widetilde{I}_{n}(\omega),&\omega\neq\pm\omega^{*}.\end{cases}$
This contamination is motivated by the case when periodic components are not suitably subtracted and are contaminated in observed time series Heyde and Dai (1996); Iacone (2010); McCloskey and Perron (2013), say, $x_{t}^{o}=x_{t}+\sqrt{8\pi z/n}\cos(At)$, $t=1,\ldots,n$ with original zero-mean time series $\{x_{t}\}_{t=1,2,\cdots,n}$. In this case, the periodogram of $x_{t}^{o}$ has approximately the form of
$I_{n}^{z,A}(\omega)=I_{n}(\omega)+z(\delta_{A}(\omega)+\delta_{-A}(\omega))+O_{P}(\sqrt{z}),$
where $\delta_{A}(\omega)=1$ if $\omega=A$ and 0 otherwise.
Observe first that the variation of the Itakura–Saito divergence with respect to an outlier in the frequency domain is written as
$D_{\mathrm{IS}}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-D_{\mathrm{IS}}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}]=\frac{z}{n}\frac{1}{S_{\theta}(\omega^{*})}-\frac{\log z}{n}+o_{P}(1)$ (5)
with $o_{P}(1)$ being uniform with respect to $z$. For larger value of $z$, this change is influenced by $\theta$, impacting on the behavior of the minimum Itakura–Saito divergence estimator. Now, putting the convex combination of $S_{\theta}$ and $\widetilde{I}^{z,\omega^{*}}_{n}$ into the second slot of the Itakura–Saito divergence mitigates the effect of $z$ as
$D_{\mathrm{IS}}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}]-D_{\mathrm{IS}}^{(n)}[\widetilde{I}_{n}\,:\,\alpha S_{\theta}+(1-\alpha)\widetilde{I}_{n}]$
$=\frac{1}{n}\frac{\alpha S_{\theta}(\omega^{*})}{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{z}{(1-\alpha)z+\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\quad-\frac{1}{n}\log(1+\widetilde{I}_{n}^{-1}(\omega^{*})z)+\frac{1}{n}\log(1+(1-\alpha)z\{\alpha S_{\theta}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})\}^{-1})$
$=o_{P}(1)$
with $o_{P}(1)$ being uniform with respect to $z$. This, together with the variational representation (3), implies that
$D_{\alpha}^{(n)}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-D_{\alpha}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}]=\frac{\alpha\log z}{n}+o_{P}(1)$ (6)
with $o_{P}(1)$ being uniform with respect to $z$. So, for large $z$, the change does not vary with respect to $\theta$. This imbues the minimum spectral Rényi divergence estimator with robustness against outliers in the frequency domain.
3 Optimization paths
This section presents a further robustness property of the minimum spectral Rényi divergence estimator. We show that (1) the optimization for the minimum spectral Rényi divergence estimator is robust against outliers in the frequency domain, whereas (2) the optimization for the minimum Itakura–Saito divergence estimator is sensitive to them.
3.1 Stable optimization path of the Spectral Rényi divergence minimization
We begin with showing that the optimization for the minimum spectral Rényi divergence estimator is robust against outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\{\gamma_{i}>0:i=1,2,\ldots\}$ of $\widehat{\theta}_{\alpha}$: for the $(m+1)$-th step,
$\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]=\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]+\gamma_{m}\,\mathcal{G}_{\alpha}(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n})$
with
$\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ $=-\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}_{n}\,:\,S_{\theta}]$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left[1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega).$ (7)
For theoretical results, we make the following assumptions. Let $\|\cdot\|$ denote the Euclidean norm.
Assumption 1.
Let $S_{\theta}\in\mathcal{S}_{\Theta}$. For each $\omega\in(0,\pi)$, the function $\theta\mapsto S_{\theta}(\omega)$ is twice continuously differentiable on $\mathrm{int}(\Theta)$, where $\mathrm{int}(\Theta)$ denotes the interior of $\Theta$. Moreover, the following conditions hold:
$\sup_{\theta\in\mathrm{int}(\Theta),\,\omega\in[-\pi,\pi]}\max\{\|\nabla_{\theta}S_{\theta}(\omega)\|,\,\|\nabla_{\theta}\log S_{\theta}(\omega)\|,\,\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\}$
$\qquad\qquad\qquad=:U_{1}<\infty;$
and
$\limsup_{n\to\infty}\left(\sup_{\theta\in\mathrm{int}(\Theta)}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)=:U_{2}<\infty,$
where $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm.
Assumption 2.
For any $z\geq 0$, $\omega^{*}\in\Omega_{n}$, and $m\in\mathbb{N}$, the $m$-th gradient descent update $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]$ lies in $\mathrm{int}(\Theta)$.
Under Assumptions 1 and 2, the subsequent theorem indicates that for any value of $\alpha\in(0,1)$, the path of the gradient descent with a prespecified step size for $\widehat{\theta}_{\alpha}$ is robust against the presence of outliers in the frequency domain.
Theorem 3 (Stability of optimization path with a fixed learning rate sequence).
Fix arbitrary step size sequence $\{\gamma_{k}>0\,:\,k=1,2,\ldots\}$. Under Assumptions 1 and 2, for any $m\in\mathbb{N}$, we have
$\lim_{n\to\infty}\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely}.$ (8)
The proof is given in Appendix B.
Remark 1.
(Constant step size)  Assumption 1 ensures that we can obtain an upper bound of the smoothness constant of $D^{(n)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ with respect to $\theta$ uniformly in $z\geq 0$: for any $z\geq 0$, we have
$\sup_{\theta\neq\theta^{\prime}}\frac{\|\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}]-\nabla_{\theta}D^{(n)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta^{\prime}}]\|}{\|\theta-\theta^{\prime}\|}$
$\leq\sup_{z\geq 0}\sup_{\theta\in\mathrm{int}(\Theta).}\|\nabla_{\theta}^{2}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|_{\mathrm{op}}$
$\leq L:=\frac{U_{1}^{2}+U_{2}}{1-\alpha}<\infty,$ (9)
where we use the following expression of the Hessian of $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$:
$\left(-\nabla_{\theta}^{2}D^{(n)}_{\alpha}[\widetilde{I}_{n}:S_{\theta}]\right)$
$=-\frac{1}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\frac{\alpha S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\qquad\qquad\frac{(1-\alpha)\widetilde{I}_{n}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\left(\nabla_{\theta}\log S_{\theta}(\omega)\right)\left(\nabla_{\theta}\log S_{\theta}(\omega)\right)^{\top}$
$\quad+\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\theta}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}\right\}\nabla^{2}_{\theta}\log S_{\theta}(\omega).$
By the standard theory of the gradient descent algorithm (c.f., Theorem $3.2$ of Nocedal and Wright (2006)), for any $z\geq 0$, the gradient along with the gradient descent update $\{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}]\}_{m=0,1,2,\ldots}$ with a constant step size $\gamma_{m}=L^{-1}$ vanishes:
$\lim_{m\to\infty}\nabla_{\theta=\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0.$
If the further assumption called the Kurdyka–Łojasiewicz inequality Łojasiewicz (1963); Kurdyka (1998) holds at all stationary points, that is, points satisfying $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]=0$ for the function $\theta\mapsto D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, the bounded gradient descent update converges to a stationary point; c.f., Theorem $3.2$ of Attouch et al. (2013). Theorem 3 implies that the finite update sequences for different $z\geq 0$ become the same and all stationary points for different $z\geq 0$ become the same as $n\to\infty$, which provides a new perspective on robust spectral analysis and optimization theory.
Remark 2.
(AR(1) model)  Here we check Assumption 1 for a first-order autoregressive model. The model spectral density is
$S_{\sigma,\rho}(\omega)=\frac{1}{2\pi}\frac{\sigma^{2}}{1-2\rho\cos(\omega)+\rho^{2}},\quad\sigma>0\quad\text{and}\quad\rho\in(-1,1).$
We reparametrize this by using
$\theta_{1}=\log\sigma\quad\text{and}\quad\theta_{2}=\,\mathrm{atanh}(\rho)$
with
$\mathrm{atanh}(\rho)=\frac{1}{2}\log\frac{1+\rho}{1-\rho}$
so as to remove constraints in the optimization. Then, the gradient of $\log S_{\theta}(\omega)$ with respect to $\theta$ is given by
$\nabla_{\theta}\log S_{\theta}(\omega)=\begin{pmatrix}2\\ 2(1-\rho^{2})\{\cos(\omega)-\rho\}/\{1-2\rho\cos(\omega)+\rho^{2}\}\end{pmatrix}$.
The Hessian of $\log S_{\theta}(\omega)$ is a bit complex but is bounded as
$\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\leq C/(1-|\rho|)^{4}$
for some absolute positive constant $C$. So, Assumption 1 holds on the interior of any compact subset $\Theta\subset(0,\infty)\times(-1,1)$.
Figure 1: Comparison of optimization paths when the length $n$ of a time series is $200$. The optimization paths of the gradient descent for the AR(1) model, with respect to $(\log\sigma,(1-\exp(-2\rho))/(1+\exp(-2\rho)))$, are presented under the spectral Rényi divergence minimization, the Itakura–Saito divergence minimization, and the time-domain $\gamma$-divergence minimization. Panel (a) shows the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $10000$ iterations (the inverted triangles), where the colors indicate different values of the contamination $z$. Panels (b) and (c) show the corresponding results for the Itakura–Saito divergence and time-domain $\gamma$-divergence minimization, respectively.
Theorem 3 establishes a connection between robustness in statistics and optimization theory, offering a novel perspective in robust statistics. While detailed numerical experiments are presented in Section 4, we briefly illustrate the implications through a simple simulation study here.
We consider contaminated observations defined as
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,n=200$
with $X_{t}$ generated from AR(1) model having $(\sigma,\rho)$. We perform 10000 iterations of gradient descent with a step size of 0.01 with respect to
$(\log\sigma,(1-\exp(-2\rho))/(1+\exp(-2\rho))).$
Figure 1 (a) illustrates optimization paths and endpoints based on $D^{(n)}_{\alpha=0.5}[I_{n}:S_{\theta}]$, starting from the initial value (gray circle). Different colors represent different contamination levels $z\in\{0,10^{1},10^{2},10^{3}\}$. The figure highlights the robustness of the optimization paths under varying contamination levels. In contrast, Figure 1 (b) shows the optimization paths and the stopping points based on the Itakura–Saito divergence $D^{(n)}_{\mathrm{IS}}[I_{n}\,:\,S_{\theta}]$ starting from the initial value (the gray circle). This figure demonstrates the non-robustness of the optimization paths of the Itakura–Saito divergence minimization under contamination. The instability observed here will be analyzed further in the subsequent subsection. Further, Figure 1 (c) shows the optimization paths and the endpoints based on the time-domain $\gamma$-divergence. Although Theorem 1 establishes a link between the time-domain $\gamma$-divergence and the spectral $\alpha$-Rényi divergence, their non-asymptotic behavior is different and the implementation of the time-domain $\gamma$-divergence minimization becomes increasingly difficult as $n$ grows, because each optimization step requires a costly log-determinant computation; see also the proof of Theorem 1 in the supplement. This comparison highlights a practical advantage of the approach using the spectral Rényi divergence.
Practically, gradient descent updates commonly use line search methods to optimize the step size dynamically. To manage this situation, we present a guiding theorem for gradient descent updates with line search. In this paper, we employ the Armijo condition for this purpose. Consider an objective function $\mathcal{L}(\theta)$ to be minimized and let $\theta^{(m)}$ be the $m$-th gradient descent update for each $m\in\mathbb{N}$. The Armijo condition with $c\in(0,1)$ selects the step size $\gamma$ in the $m$-th update $\theta^{(m)}$ so as to satisfy
$\mathcal{L}\left(\theta^{(m-1)}-\gamma\nabla_{\theta=\theta^{(m-1)}}\mathcal{L}(\theta)\right)\leq\mathcal{L}\left(\theta^{(m-1)}\right)-c\gamma\left\|\nabla_{\theta=\theta^{(m-1)}}\mathcal{L}(\theta)\right\|^{2},$
which ensures the monotone decrease of the objective function along with the gradient descent update. We denote by $\mathcal{A}[c,\theta,\mathcal{L}]$ the Armijo condition with $c\in(0,1)$ for an objective function $\mathcal{L}(\theta)$ in the update from $\theta$.
Proposition 2 (Stability of line search).
Fix $c\in(0,1)$, $\kappa>0$, and $\underline{\gamma}>0$. Fix also the point $\theta\in\Theta$ such that
$\left\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\right\|>\kappa.$
[…]
The proof is given in Appendix C. It employs the descent lemma (c.f., Bertsekas (1999)) together with the comparison of gradients.
Remark 3.
(Optimization path with the Armijo condition)  Let $\{\gamma_{m}[\widetilde{I}_{n}]:m=1,2,\ldots\}$ and $\{\gamma_{m}[\widetilde{I}^{z,\omega^{*}}_{n}]:m=1,2,\ldots\}$ the step size sequences for $D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]$ and $D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$, respectively. From Proposition 2, we can assume
$\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}|\gamma_{m}[\widetilde{I}_{n}]-\gamma_{m}[\widetilde{I}_{n}^{z,\omega^{*}}]|\to 0\quad\text{as }\quad n\to\infty.$
Under this additional assumption and if we have
$\inf_{k=0,\ldots,m-1}\|\nabla_{\theta=\widehat{\theta}_{\alpha}^{(k)}[\widetilde{I}_{n}]}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\|>0,$
we can have for any $m\in\mathbb{N}$,
$\lim_{n\to\infty}\sup_{z\geq 0,\omega^{*}\in\Omega_{n}}\left\|\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}_{\alpha}^{(m)}[\widetilde{I}_{n}]\right\|=0\text{ almost surely},$
even when we utilize the line search.
3.2 Unstable optimization path of the Itakura–Saito divergence minimization
We next show that the optimization for the minimum Itakura–Saito divergence estimator is sensitive to outliers in the frequency domain. Consider the gradient descent update with a sequence of learning rates $\{\gamma_{i}>0:i=1,2,\ldots\}$ of $\widehat{\theta}_{\mathrm{IS}}$: for the $(m+1)$-th step,
$\widehat{\theta}_{\mathrm{IS}}^{(m+1)}[\widetilde{I}_{n}]=\widehat{\theta}_{\mathrm{IS}}^{(m)}[\widetilde{I}_{n}]+\gamma_{m}\,\mathcal{G}_{\mathrm{IS}}(\widehat{\theta}_{\mathrm{IS}}^{(m)}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n})$
with
$\mathcal{G}_{\mathrm{IS}}(\theta\,;\,\widetilde{I}_{n})=\frac{1}{n}\sum_{\omega\in\Omega_{n}}\left[\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}-1\right]\nabla_{\theta}\log S_{\theta}(\omega).$
The subsequent proposition suggests that both the initial update and the convergent point for $\widehat{\theta}_{\mathrm{IS}}$ are highly sensitive to outliers in the frequency domain.
Proposition 3.
Fix $n\in\mathbb{N}$. Assume that for $\omega^{*}\in\Omega_{n}$, the inequality
$\inf_{\theta\in\mathrm{int}(\Theta)}\|(1/S_{\theta}(\omega^{*}))\nabla_{\theta}\log S_{\theta}(\omega^{*})\|>0$
holds and $\theta^{(0)}$ lies in $\mathrm{int}(\Theta)$. Then we have
$\sup_{z>0}\left\|\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]\right\|=\infty\,\,\text{almost surely}.$ (10)
Further, assume that for any $z>0$ and $\omega^{*}\in\Omega_{n}$, the gradient descent sequence $\{\widehat{\theta}^{(m)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}]\}_{m=1,2\ldots}$ lies in $\mathrm{int}(\Theta)$. Then, even when the gradient descent sequence with $\widetilde{I}^{z,\omega^{*}}_{n}$ has a subsequence converging to a stationary point $\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]$ for $D_{\mathrm{IS}}^{(n)}[\,\widetilde{I}^{z,\omega^{*}}_{n}\,:\,S_{\theta}\,]$, i.e.
$\nabla_{\theta=\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}D_{\mathrm{IS}}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}]=0,$
this stationary point $\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]$ is not a stationary point for $D_{\mathrm{IS}}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]$:
$\sup_{z>0}\left\|\nabla_{\theta=\widehat{\theta}^{(\infty)}[\widetilde{I}_{n}^{z,\omega^{*}}]}D_{\mathrm{IS}}^{(n)}[\,\widetilde{I}_{n}\,:\,S_{\theta}\,]\right\|=\infty\,\,\text{almost surely}.$
The proof is given in Appendix D.
4 Simulation studies
This section presents two numerical studies using the AR(2) model and the Brune spectral model with attenuation that is motivated by seismological studies Yoshimitsu et al. (2023).
4.1 AR(2) model
Table 1: The mean values of estimates with standard deviations for the AR(2) model without any trend. The values closest to the true values are underlined.
$\widehat{\sigma}$ $\widehat{\varphi_{1}}$ $\widehat{\varphi_{2}}$
Rényi with $\alpha=0.50$ 0.83 ($\pm$ 0.06) 0.57 ($\pm$ 0.64) 0.14 ($\pm$ 0.65)
Rényi with $\alpha=0.75$ 0.95 ($\pm$ 0.10) 0.84 ($\pm$ 0.50) -0.17 ($\pm$ 0.44)
Rényi with $\alpha=0.90$ 1.02 ($\pm$ 0.12) 1.01 ($\pm$ 0.40) -0.28 ($\pm$ 0.38)
IS with the periodogram 1.13 ($\pm$ 0.18) 1.09 ($\pm$ 0.48) -0.32 ($\pm$ 0.42)
IS with the smoothed periodogram 1.14 ($\pm$ 0.19) 1.11 ($\pm$ 0.51) -0.34 ($\pm$ 0.44)
Table 2: The mean values of estimates with standard deviations for the AR(2) model with a trigonometric trend. The values closest to the true values are underlined.
$\widehat{\sigma}$ $\widehat{\varphi_{1}}$ $\widehat{\varphi_{2}}$
Rényi with $\alpha=0.50$ 0.85 ($\pm$ 0.05) 0.53 ($\pm$ 0.68) 0.20 ($\pm$ 0.74)
Rényi with $\alpha=0.75$ 0.99 ($\pm$ 0.12) 0.81 ($\pm$ 0.58) -0.14 ($\pm$ 0.51)
Rényi with $\alpha=0.90$ 1.07 ($\pm$ 0.14) 1.02 ($\pm$ 0.46) -0.26 ($\pm$ 0.42)
IS with the periodogram 2.35 ($\pm$ 0.09) 0.02 ($\pm$ 0.05) 0.98 ($\pm$ 0.05)
IS with the smoothed periodogram 2.35 ($\pm$ 0.10) 0.03 ($\pm$ 0.06) 0.97 ($\pm$ 0.06)
First, we check the performance using the AR(2) model, where the spectral density is given by
$S^{\mathrm{AR2}}_{\theta}(\omega):=\sigma^{2}\Big/\left|1-\varphi_{1}e^{-\sqrt{-1}\omega}-\varphi_{2}e^{-2\sqrt{-1}\omega}\right|^{2},\;\;\theta=(\sigma,\varphi_{1},\varphi_{2}).$
We use the true value $\theta^{\ast}:=(\sigma^{\ast},\varphi_{1}^{\ast},\varphi_{2}^{\ast})=(1,0.9,-0.2)$. We use three different initial values: $\theta^{(0),1}=(0.1,0,0)$, $\theta^{(0),2}=(1,0,0)$, and $\theta^{(0),3}=(10,0,0)$. The optimization is constrained to the stationary region of the AR(2) process; see Section E for details. We generate a time series $\mathbf{x}_{n=2^{10}}$ with its spectral density $S_{\theta^{*}}^{\mathrm{AR2}}$ 1000 times, and if we consider the contamination in the frequency domain, we add trigonometric trends to the time series as
$x_{t}^{o}=x_{t}+z_{1}\sqrt{2\pi/n}\sin(A_{1}t)+z_{2}\sqrt{2\pi/n}\sin(A_{2}t),\,\,t=1,\ldots,n$
with $z_{1}=z_{2}=20$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$.
We compare five estimation methods: (1,2,3) the minimum spectral Rényi divergence estimators with
$\alpha\in\{0.5,0.75,0.9\}$
using the periodogram ($I_{n}$); (4) the minimum Itakura–Saito divergence estimator employing the periodogram ($I_{n}$); and (5) the minimum Itakura–Saito divergence estimator with the smoothed periodogram ($I_{n}^{S}$).
For each divergence minimization problem, the minimization is conducted via the gradient descent algorithm with a fixed learning rate $\gamma=0.005$. We stop the optimization when the optimization step reaches 10000 or when the Euclidean norm of the gradient of each divergence evaluated at each updated estimate is less than $10^{-3}$. For smoothing the periodogram, we apply the modified Daniell smoothers twice, where the length of the first time is $3$ and the length of the second time is $5$.
Tables 1 and 2 summarize the estimation results in the above setup. Tables 1 and 2 suggest that the minimizers of the IS divergence with the periodogram/smoothed periodogram are very sensitive to the trigonometric trends of the observed time series (i.e. the outliers in the frequency domain). On the other hand, the minimizers of the spectral Rényi divergence give stably estimated values and their performance would be substantially better than those of the IS divergence with the periodogram/smoothed periodogram. Thus the estimated values using the spectral Rényi divergence are robust against the outliers in the frequency domain.
Figure 2: Spectral densities with the estimates plugged in for the AR(2) model without any trend. Each panel shows the results for different realizations of the periodogram. The gray curve is the periodogram. The green curve is the smoothed periodogram. The true spectral density is colored in black. The spectral density based on the spectral Rényi divergence is colored in red. The spectral density based on the Itakura–Saito divergence is colored in blue. We set $\theta^{(0),1}$ to the initial value. The initial value $\theta^{(0),\mathrm{best}}$ represents the point that minimizes the objective functions over a grid of the parameter space.
Figure 3: Spectral densities with the estimates plugged in for the AR(2) model with the trend. Each panel shows the results for different realizations of the periodogram. The legend is the same as in Figure 2. The red dot lines represent the points of the outliers. We set $\theta^{(0),1}$ to the initial value.
Figure 4: Histograms of the numbers of iterations to converge in the AR(2) model, where the initial value is $\theta^{(0),2}$. We conducted each simulation 100 times and plotted the histogram. (a) The Itakura–Saito divergence minimization without any trend. (b) The Itakura–Saito divergence minimization with a trend. (c) The $\alpha=0.9$-Rényi divergence minimization without any trend. (d) The $\alpha=0.9$-Rényi divergence minimization with a trend.
Figures 2 and 3 display the estimated spectral densities. In each figure, the solid gray curve represents the periodogram $I_{n}(\omega)$; the solid green curve represents the smoothed periodogram $I_{n}^{S}(\omega)$; the solid black curve represents the spectral density $S^{\mathrm{AR2}}_{\theta}(\omega)$ with the true value $\theta^{*}$ plugged in; the solid blue curve represents the spectral density with the minimum Itakura–Saito divergence estimate plugged in; the solid red curve represents the spectral density with the minimum spectral Rényi divergence ($\alpha=0.9$) estimate plugged in. The dashed curves represent the estimates starting from the initial values that minimize the objective functions, respectively. Here the periodogram is used in these figures. Figures 2 and 3 show that the spectral density $S^{\mathrm{AR2}}_{\widehat{\theta}_{n}}$ with the minimum spectral Rényi divergence ($\alpha=0.9$) (resp. IS divergence) estimate $\widehat{\theta}_{n}$ plugged in fits (resp. does not fit) well to the spectral density $S^{\mathrm{AR2}}_{\theta^{*}}$ with the true value $\theta^{*}$ plugged in, which implies that the minimizers of the IS divergence with the periodogram/smoothed periodogram are very sensitive to the outliers in the frequency domain, but the spectral Rényi divergence is robust against them. These figures also indicate that the Itakura–Saito divergence estimate is highly sensitive to realizations of the periodogram even without any contamination. The grid search for the initial values does not help improve the performance of the Itakura–Saito divergence estimate. In contrast, the spectral Rényi estimate is generally less sensitive to realizations of the periodogram.
Further, Figure 4 displays the histograms of the numbers of iterations required for convergence when the initial value is $\theta^{(0),2}$. For the Rényi-divergence minimization, the vast majority of trials converge within a moderate number of iterations, both with and without outliers. In contrast, for the Itakura–Saito divergence minimization, the convergence behavior deteriorates markedly in the presence of outliers, where essentially all trials reach the maximum number of iterations without convergence. These findings are consistent with the optimization trajectories illustrated in Figure 1.
4.2 The Brune spectral model with attenuation
Next, consider the Brune spectral model Brune (1970) with attenuation Aki and Richards (1980) for velocity waveforms:
$S^{\mathrm{BA}}_{\theta}(\omega):=\omega^{2}\left[\frac{\sigma^{2}}{\{1+(\omega/\omega_{c})^{2}\}^{2}}\right]\exp(-\omega/Q),\;\;\theta=(\sigma,\omega_{c},Q).$
The Brune spectral model $\sigma^{2}/\{1+(\omega/\omega_{c})^{2}\}^{2}$ for displacement waveforms is the spectral density model of Matérn process Matérn (1960) with the spectral decay rate fixed to 2 and often used in the earthquake source estimation Calderoni and Abercrombie (2023); Yoshimitsu et al. (2023). The attenuation factor $\exp(-\omega/Q)$ represents anelasticity of the medium through which the seismic wave propagates. The multiplicative factor $\omega^{2}$ comes from the time derivative that converts a displacement waveform to a velocity waveform.
Figure 5: Spectral densities with the estimates plugged in for the Brune model without any trend. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 6: Spectral densities with the estimates plugged in for the Brune model with the trigonometric trends. The red dashed lines denote the frequencies at which the outliers are injected. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Table 3: Mean values of biases with standard deviations for the Brune model without any trend. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),1}$ -0.03 ($\pm$0.05) -0.12 ($\pm$0.07) 0.30 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.17 ($\pm$0.06) 0.54 ($\pm$0.22)
R ($\alpha=0.90$) $\theta^{(0),1}$ 0.01 ($\pm$0.06) -0.20 ($\pm$0.06) 0.90 ($\pm$0.33)
IS with $I_{n}$ $\theta^{(0),1}$ 42.6 ($\pm$2.66) 893 ($\pm$50.5) 48.2 ($\pm$2.99)
IS with $I_{n}^{S}$ $\theta^{(0),1}$ 3372 ($\pm$5318) 67386 ($\pm$106260) 5259 ($\pm$8327)
R ($\alpha=0.50$) $\theta^{(0),2}$ -0.03 ($\pm$0.05) -0.10 ($\pm$0.07) 0.23 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),2}$ 0.002 ($\pm$0.06) -0.17 ($\pm$0.07) 0.50 ($\pm$0.21)
R ($\alpha=0.90$) $\theta^{(0),2}$ 0.01 ($\pm$0.06) -0.23 ($\pm$0.05) 1.05 ($\pm$0.34)
IS with $I_{n}$ $\theta^{(0),2}$ 0.06 ($\pm$0.10) -0.04 ($\pm$0.13) 0.06 ($\pm$0.17)
IS with $I_{n}^{S}$ $\theta^{(0),2}$ -0.29 ($\pm$0.30) 1.95 ($\pm$1.89) 1.86 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$ -0.10 ($\pm$0.04) 0.51 ($\pm$0.13) -0.30 ($\pm$0.06)
R ($\alpha=0.75$) $\theta^{(0),3}$ -0.05 ($\pm$0.06) 0.08 ($\pm$0.19) 0.05 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),3}$ -0.007 ($\pm$0.06) -0.15 ($\pm$0.13) 0.72 ($\pm$0.35)
IS with $I_{n}$ $\theta^{(0),3}$ 0.10 ($\pm$0.07) 0.42 ($\pm$0.26) -0.28 ($\pm$0.13)
IS with $I_{n}^{S}$ $\theta^{(0),3}$ -0.29 ($\pm$0.15) 2.09 ($\pm$1.68) 1.50 ($\pm$0.72)
Table 4: Mean values of biases with standard deviations for the Brune model with the trigonometric trends. Values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),1}$ 0.22 ($\pm$0.07) -0.22 ($\pm$0.05) 0.27 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),1}$ 0.53 ($\pm$0.08) -0.34 ($\pm$0.04) 0.47 ($\pm$0.19)
R ($\alpha=0.90$) $\theta^{(0),1}$ 1.45 ($\pm$0.10) -0.48 ($\pm$0.03) 0.74 ($\pm$0.31)
IS with $I_{n}$ $\theta^{(0),1}$ 75.8 ($\pm$2.85) 1524 ($\pm$55.86) 60.0 ($\pm$3.04)
IS with $I_{n}^{S}$ $\theta^{(0),1}$ 3405 ($\pm$5318) 68019 ($\pm$106263) 5271 ($\pm$8327)
R ($\alpha=0.50$) $\theta^{(0),2}$ 0.21 ($\pm$0.07) -0.20 ($\pm$0.06) 0.23 ($\pm$0.14)
R ($\alpha=0.75$) $\theta^{(0),2}$ 0.53 ($\pm$0.08) -0.33 ($\pm$0.04) 0.45 ($\pm$0.19)
R ($\alpha=0.90$) $\theta^{(0),2}$ 1.48 ($\pm$0.10) -0.49 ($\pm$0.03) 0.79 ($\pm$0.33)
IS with $I_{n}$ $\theta^{(0),2}$ 3.92 ($\pm$0.07) -0.35 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),2}$ 2.83 ($\pm$0.97) -0.09 ($\pm$0.97) 2.30 ($\pm$0.98)
R ($\alpha=0.50$) $\theta^{(0),3}$ 0.08 ($\pm$0.06) 0.24 ($\pm$0.15) -0.25 ($\pm$0.10)
R ($\alpha=0.75$) $\theta^{(0),3}$ 0.46 ($\pm$0.08) -0.27 ($\pm$0.07) 0.26 ($\pm$0.16)
R ($\alpha=0.90$) $\theta^{(0),3}$ 1.40 ($\pm$0.09) -0.47 ($\pm$0.03) 0.69 ($\pm$0.29)
IS with $I_{n}$ $\theta^{(0),3}$ 3.90 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),3}$ 2.80 ($\pm$0.95) -0.13 ($\pm$0.59) 2.01 ($\pm$0.86)
For the experiments, we use the true value $\theta^{\ast}:=(\sigma^{\ast},\omega_{c}^{\ast},Q^{\ast})=(1,1,1)$. We generate a time series $\mathbf{x}_{n=2^{10}}$ with its spectral density $S_{\theta^{*}}^{\mathrm{BA}}$ 100 times, and if we consider outliers in the frequency domain, we add two trigonometric trends to the time series as
$x_{t}^{o}=x_{t}+\sqrt{8\pi z_{1}/n}\sin(A_{1}t)+\sqrt{8\pi z_{2}/n}\sin(A_{2}t),\,\,t=1,\ldots,n$
with $z_{1}=z_{2}=100$, $A_{1}=\pi/4$ and $A_{2}=\pi/8$. We report the results based on different values of $z_{1}=z_{2}$ in Appendix F. We use three different initial values: $\theta^{(0),1}=(1,0.1,1)$, $\theta^{(0),2}=(1,1,1)$, and $\theta^{(0),3}=(1,2,1)$. Setting $\sigma^{*}$ and $Q^{*}$ as initial values does not impact on the behavior of the minimum spectral Rényi divergence estimates; we report the results based on different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$ in Appendix F.
Figures 5 and 6 display the estimated spectral densities. In each figure, the following representations are used:
  • The solid gray curve represents the periodogram $I_{n}$;
  • the solid black curve denotes the true spectral density $S^{\mathrm{BA}}_{\theta^{*}}$;
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged in;
  • the solid red, salmon pink, and green curves display the spectral densities with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged in, respectively.
For any initial value and regardless of the existence of the trend, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence aligns well with the true spectral density. The minimum spectral Rényi divergence estimate with any $\alpha\in\{0.5,0.75,0.9\}$ remains stable with respect to the choice of the initial value. Conversely, the minimum Itakura–Saito divergence estimate shows sensitivity to this choice.
Tables 3 and 4 summarize the estimation results. Comparing the results based on different initial values, we find that the minimum Itakura–Saito divergence estimates are sensitive to the initial value of the optimization regardless of the existence of the smoother, while the spectral Rényi divergence yields estimation results stable to the choice of the initial value. The minimum Itakura–Saito divergence estimate with the smoothed periodogram performs much worse than that with the periodogram. In the presence of outliers in the frequency domain, the minimum Rényi divergence estimate with $\alpha=0.5$ performs the best regardless of the choice of the initial value in this example.
5 Discussions
Finally, we discuss three important directions for future research.
The first point is the choice of $\alpha$. The value of $\alpha$ controls the trade-off between the efficiency without the presence of outliers and the robustness even for larger trends. For an objective selection of $\alpha$, constructing an information criterion is one possible direction, but requires sufficient discussions and experiments. In practice, monitoring the behaviors of the minimum spectral Rényi divergence estimates for several values of $\alpha$ (say, $\alpha=0.5,0.75,0.9$) would be a workaround.
The second point is the choice of optimization schemes. In this paper, we employed a simple gradient descent scheme only as a baseline illustrative method, and we do not claim that naive gradient descent is the most appropriate choice in practice. As is well known, the iteration number, tolerance, and step-size selection in naive gradient descent can be highly sensitive and not objectively determined. There are at least two ways to avoid these issues:
  1. 1.
    Use of line-search methods. One can incorporate standard line-search strategies such as Wolfe or Armijo conditions (see Proposition 2 and Remark 3), or alternatively employ recent line-search-free first-order methods Malitsky and Mishchenko (2020); Yagishita and Ito (2025), which often provide more stable convergence behavior without requiring excessively large iteration limits.
  2. 2.
    Use of quasi-Newton methods. Quasi-Newton algorithms such as the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm or L-BFGS Liu and Nocedal (1989) are typically more efficient in terms of optimization and are widely preferred by practitioners. Indeed, in a large-scale application of spectral Rényi divergence minimization to real geophysical data Kano et al. (2025), we adopted the L-BFGS method precisely to avoid the practical difficulties associated with naive first-order methods such as gradient descent.
Constructing a unified theory that accommodates line-search-free first-order methods and quasi-Newton methods would be an important research direction.
The third one is the impact of the choice of an initial value. Numerical experiments presented in Section 4 suggest that the minimum spectral Rényi divergence estimates are robust to the choice of an initial value, while the minimum Itakura–Saito divergence estimate is sensitive to it. These robustness and sensitivity are confirmed in the other examples such as autoregressive models. The structure of the gradient $\mathcal{G}_{\alpha}(\theta\,;\,\widetilde{I}_{n})$ in the comparison with that of $\mathcal{G}_{\mathrm{IS}}(\theta\,;\,\widetilde{I}_{n})$ can give a clue to this, and so detailed analysis of these gradient landscapes would be one of the interesting research directions.
6 Acknowledgements
The authors would like to thank the Associate Editor and the referee for their constructive comments, which improved the quality of the paper. The authors would like to thank Akifumi Okuno, Mirai Tanaka, Kei Kobayashi, Yuta Koike, Tomoyuki Higuchi, Hironori Fujisawa, Junho Yang, Masayuki Kano, and Shotaro Yagishita for their comments. This work is supported by JSPS KAKENHI (19K20222, 21H05205, 21K12067, 23K11024), MEXT (JPJ010217), and “Strategic Research Projects” grant (2022-SRP-13) from ROIS (Research Organization of Information and Systems). The simulation code is available at https://github.com/t-tetsuya/Spectral_Renyi.git.
Supplement to “On robustness of spectral Rényi divergence” by Tetsuya Takabatake and Keisuke Yano
This supplement presents
  • Proof of Theorem 1 in Appendix A;
  • Proof of Theorem 3 in Appendix B;
  • Proof of Proposition 2 in Appendix C;
  • Proof of Proposition 3 in Appendix D;
  • Additional comparison of optimization paths in Appendix E; and
  • Additional simulation studies in Appendix F.
Appendix A Proof of Theorem 1
We first introduce the notation used in the proof. For each spectral density function $S$ of a stationary time series $(X_{t})_{t\in\mathbb{Z}}$, we write $\Sigma_{n}(S)$ as the $n\times n$-Toeplitz matrix whose $(s,t)$th element is given by
$R(t-s)=\int_{-\pi}^{\pi}e^{\sqrt{-1}(t-s)\omega}S(\omega)\,\mathrm{d}\omega,\ \ s,t=1,\ldots,n,$
which corresponds to the covariance matrix of $(X_{1},X_{2},\cdots,X_{n})^{\top}$. Then, the probability density function of $(X_{1},\cdots,X_{n})^{\top}$ is expressed by
$p_{S}(\mathbf{x})=\frac{1}{\sqrt{(2\pi)^{n}\det[\Sigma_{n}(S)]}}\exp\left(-\frac{1}{2}\mathbf{x}^{\top}\Sigma_{n}(S)^{-1}\mathbf{x}\right),\ \ \mathbf{x}\in\mathbb{R}^{n}.$ (11)
Let $\gamma\in(0,\infty)$. First introduce the $\gamma$-cross entropy $d_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,$] between two probability density functions $p_{S}$ and $p_{\widetilde{S}}$, which are not necessarily Gaussian probability density functions, defined by
$d_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,]:=-\frac{1}{\gamma}\log\int_{\mathbb{R}^{n}}p_{S}(\mathbf{x})p_{\widetilde{S}}(\mathbf{x})^{\gamma}\,\mathrm{d}\mathbf{x}+\frac{1}{1+\gamma}\log\int_{\mathbb{R}^{n}}p_{\widetilde{S}}(\mathbf{x})^{1+\gamma}\,\mathrm{d}\mathbf{x}.$
Then the $\gamma$-divergence can be written as $G_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,]=d_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,]-d_{\gamma}[\,p_{S}\,:\,p_{S}\,]$.
[…]
$\int_{\mathbb{R}^{n}}p_{S}(\mathbf{x})p_{\widetilde{S}}(\mathbf{x})^{\gamma}\,\mathrm{d}\mathbf{x}$ $=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
$\qquad\quad\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}\mathbf{x}^{\top}[\Sigma_{n}(S)^{-1}+\gamma\Sigma_{n}(\widetilde{S})^{-1}]\mathbf{x}\right)\,\mathrm{d}\mathbf{x}$
$=\left\{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}\right\}^{-\frac{1}{2}}$
$\qquad\quad\int_{\mathbb{R}^{n}}\exp\left(-\frac{1}{2}\mathbf{x}^{\top}\Sigma_{n}(S)^{-1}[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]\Sigma_{n}(\widetilde{S})^{-1}\mathbf{x}\right)\,\mathrm{d}\mathbf{x}$
$=\left\{\frac{\pi^{n}\det[\Sigma_{n}(\widetilde{S})(\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S))^{-1}\Sigma_{n}(S)]}{\pi^{n(1+\gamma)}\det[\Sigma_{n}(S)]\det[\Sigma_{n}(\widetilde{S})]^{\gamma}}\right\}^{\frac{1}{2}}$
$=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]^{-1}\right\}^{\frac{1}{2}}.$
Similarly, we also obtain
$\int_{\mathbb{R}^{n}}p_{\widetilde{S}}(\mathbf{x})^{1+\gamma}\,\mathrm{d}\mathbf{x}$ $=\left\{\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[(1+\gamma)\Sigma_{n}(\widetilde{S})]^{-1}\right\}^{\frac{1}{2}}$
$=\left\{\pi^{-n\gamma}(1+\gamma)^{-n}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right\}^{\frac{1}{2}}.$
Combining these yields
$2d_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,$]= $-\frac{1}{\gamma}\log\left[\pi^{-n\gamma}\det[\Sigma_{n}(\widetilde{S})]^{1-\gamma}\det[\Sigma_{n}(\widetilde{S})+\gamma\Sigma_{n}(S)]^{-1}\right]$
$\qquad+\frac{1}{1+\gamma}\log\left[\pi^{-n\gamma}(1+\gamma)^{-n}\det[\Sigma_{n}(\widetilde{S})]^{-\gamma}\right]$
$=$ $n\log(\pi)-\frac{n}{1+\gamma}\log[(1+\gamma)\pi^{\gamma}]$
[…]
Putting $S$ into $\widetilde{S}$, we obtain
$2d_{\gamma}[\,p_{S}\,:\,p_{S}\,]$ $=n\log(\pi)-\frac{n}{1+\gamma}\log[(1+\gamma)\pi^{\gamma}]+\frac{1}{\gamma}\log\det[(1+\gamma)\Sigma_{n}(S)]$
$\qquad\qquad\qquad\qquad\qquad-\frac{1}{\gamma(1+\gamma)}\log\det[\Sigma_{n}(S)]$
$=n\log(\pi)-\frac{n}{1+\gamma}\log[(1+\gamma)\pi^{\gamma}]+\frac{n}{\gamma}\log(1+\gamma)+\frac{1}{1+\gamma}\log\det[\Sigma_{n}(S)].$
Thus we get
[…]
$\qquad\qquad\qquad\qquad\qquad\qquad-\log(1+\gamma)^{n}-\frac{\gamma}{1+\gamma}\log\det[\Sigma_{n}(S)]\Bigg]$
[…]
where $\bar{S}^{(\gamma)}(\omega):=\frac{\gamma}{1+\gamma}S(\omega)+\frac{1}{1+\gamma}\widetilde{S}(\omega)$. Using Szegö’s limit theorem Szegö (1920); Kolmogorov (1941), we conclude
$\lim_{n\to\infty}\frac{2}{n}G_{\gamma}[\,p_{S}\,:\,p_{\widetilde{S}}\,$] $=\frac{1}{2\pi\gamma}\int_{-\pi}^{\pi}\left[\log{\bar{S}^{(\gamma)}(\omega)}-\frac{\gamma}{1+\gamma}\log{S(\omega)}-\frac{1}{1+\gamma}\log{\widetilde{S}(\omega)}\right]\,\mathrm{d}\omega$
$=D_{\alpha}[\,S\,:\,\widetilde{S}\,$],
where $\alpha:=(1+\gamma)^{-1}$. This completes the proof.
Appendix B Proof of Theorem 3
We prove Theorem 3 by the mathematical induction.
Step 1: Consider $m=1$. Observe
$\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]$
$=\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\right\}-\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]-\theta^{(0)}\right\}$
$=\frac{\gamma_{1}\alpha}{(1-\alpha)n}\Bigg(\frac{S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\qquad\qquad\qquad\qquad-\frac{S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)(\widetilde{I}_{n}(\omega^{*})+z)}\Bigg)\nabla_{\theta^{(0)}}\log S_{\theta}(\omega^{*})$
$=\frac{\gamma_{1}}{(1-\alpha)n}\frac{\alpha S_{\theta^{(0)}}(\omega^{*})}{\alpha S_{\theta^{(0)}}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{(1-\alpha)z\nabla_{\theta^{(0)}}S_{\theta}(\omega^{*})}{(1-\alpha)z+(1-\alpha)\widetilde{I}_{n}(\omega^{*})+\alpha S_{\theta^{(0)}}(\omega^{*})}.$
This yields
$\left\|\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]\right\|\leq\frac{1}{n}\frac{\gamma_{1}}{1-\alpha}\|\nabla_{\theta^{(0)}}S_{\theta}(\omega^{*})\|,$ (12)
which proves the assertion for $m=1$.
Step 2: Let $m\in\mathbb{N}>1$. Assume that for any $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for $n\geq N$, we have
$\|\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\|\leq\varepsilon.$
By the cancelling technique and by the triangle inequality, we get
$\left\|\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]\right\|$ $=\Bigg\|\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]\right\}$
$\qquad-\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\right\}-\left\{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\right\}\Bigg\|$
$\leq\gamma_{m}\left\|\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)\right\|+\varepsilon.$
Observe that by applying the Taylor theorem to $\nabla_{\theta}\log S_{\theta}(\omega)$ and letting $t$ be some point on some line connecting $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]$ to $\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]$, we get
$\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega)$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{t}^{2}\log S_{\theta}(\omega)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}])$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left\{1-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\right\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\mathrm{int}(\Theta)}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
where $O(1)$ in the rightmost side implies a vector whose norm is $O(1)$. This yields
$\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)$
$=\frac{\alpha}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\Bigg\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\qquad\qquad-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\Bigg\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\mathrm{int}(\Theta)}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)$
$=\frac{\alpha}{(1-\alpha)n}(T_{1}+T_{2})+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\mathrm{int}(\Theta)}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
where
$T_{1}$ $:=\sum_{\omega\in\Omega_{n}}\Bigg\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}_{n}(\omega)}$
$\qquad\qquad\quad-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\Bigg\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega),$
$T_{2}$ $:=\sum_{\omega\in\Omega_{n}}\Bigg\{\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}$
$\qquad\qquad\quad-\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\Bigg\}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega).$
Here we have
$T_{1}$ $=\frac{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}\frac{(1-\alpha)z}{(1-\alpha)z+\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega^{*})+(1-\alpha)\widetilde{I}_{n}(\omega^{*})}$
$\qquad\frac{\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega^{*})}{\alpha}$
$=O\left(\frac{1}{\alpha}\sup_{\theta,\,\omega}\|\nabla_{\theta}\log S_{\theta}(\omega)\|\right)$
and
$T_{2}$ $=\sum_{\omega\in\Omega_{n}}\frac{(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}$
$\qquad\qquad\frac{S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)-S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega)}{\alpha S_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}(\omega)+(1-\alpha)\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}\nabla_{\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]}\log S_{\theta}(\omega)$
$=O\left(\frac{n\varepsilon}{\alpha}\sup_{\theta,\theta^{\prime},\omega}\|\nabla_{\theta^{\prime}}S_{\theta}(\omega)\|\|(1/S_{\theta}(\omega))\nabla_{\theta}\log S_{\theta}(\omega)\|\right).$
Then we obtain
$\left\|\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]\,;\,\widetilde{I}^{z,\omega^{*}}_{n}\right)-\mathcal{G}_{\alpha}\left(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}]\,;\,\widetilde{I}_{n}\right)\right\|$
$=\frac{1}{(1-\alpha)n}O(U_{1})+\varepsilon\Bigg\{\frac{1}{(1-\alpha)}O(U_{1}^{2})+\frac{2}{(1-\alpha)}O(U_{2})\Bigg\}$
and thus
$\left\|\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]\right\|$
$\leq\frac{\gamma_{m}}{(1-\alpha)n}O(U_{1})+\varepsilon\Bigg\{1+\frac{\gamma_{m}}{(1-\alpha)}O(U_{1}^{2})+\frac{2\gamma_{m}}{(1-\alpha)}O(U_{2})\Bigg\}.$
This implies that the assertion holds for $m+1$ and by the mathematical induction, we get the conclusion. ∎
Appendix C Proof of Proposition 2
Let $\gamma$ be the step size satisfying the Armijo condition $\mathcal{A}[c\,,\,\theta\,,\,D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]]$ and $\gamma\leq 2(1-c)/L$. Let $\mathcal{P}_{\gamma}[\theta]$ be
$\mathcal{P}_{\gamma}[\theta]:=\theta-\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}\,:\,S_{\theta}].$
We begin with utilizing the descent lemma (c.f., Bertsekas (1999)):
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]$ $\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]+\langle\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\,,\,\mathcal{P}_{\gamma}[\theta]-\theta\rangle$
$\quad+\frac{L}{2}\|\mathcal{P}_{\gamma}[\theta]-\theta\|^{2},$ (13)
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner-product. From equation (12), we have
$\left|\left\langle v\,,\,(\mathcal{P}_{\gamma}[\theta]-\theta)\right\rangle-\left\langle v\,,\,\left(-\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\right)\right\rangle\right|$ $\leq\frac{1}{n}\frac{\gamma}{1-\alpha}\|v\|\|\nabla_{\theta}S_{\theta}(\omega^{*})\|$
$\leq\frac{1}{n}\frac{2U_{1}}{L(1-\alpha)}\|v\|,$
where $U_{1}$ is defined in Assumption 1 and $L$ is defined in (9). Putting the quantity $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$ into $v$ in the above inequality, we get
$\langle\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\,,\,\mathcal{P}_{\gamma}[\theta]-\theta\rangle\leq-\gamma\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\frac{1}{n}\frac{2U_{1}^{2}}{L(1-\alpha)^{2}},$ (14)
where we use the inequality
$\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|\leq\frac{\alpha U_{1}}{(1-\alpha)n}\sum_{\omega\in\Omega_{n}}\left|\frac{(\alpha-1)S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}^{z,\omega^{*}}(\omega)}{\alpha S_{\theta}(\omega)+(1-\alpha)\widetilde{I}_{n}^{z,\omega^{*}}(\omega)}\right|\leq\frac{U_{1}}{1-\alpha};$
see (7) for the expression of $\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]$. Equation (12) also gives
$\left\|\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]-\gamma\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\right\|\leq\frac{\gamma U_{1}}{n(1-\alpha)}\leq\frac{2U_{1}}{nL(1-\alpha)}$
so that we obtain
$\|\mathcal{P}_{\gamma}[\theta]-\theta\|^{2}\leq\gamma^{2}\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\left(\frac{2}{n}+\frac{1}{n^{2}}\right)\frac{4U_{1}^{2}}{L^{2}(1-\alpha)^{2}}.$ (15)
Together with (13), equations (14) and (15) imply
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]-\left(\gamma-\frac{L}{2}\gamma^{2}\right)\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\varepsilon_{n},$
where $\varepsilon_{n}$ is given by
$\varepsilon_{n}:=\left(\frac{3}{n}+\frac{1}{n^{2}}\right)\frac{2U_{1}^{2}}{L(1-\alpha)^{2}}.$
For $\gamma$ satisfying $1-(L/2)\gamma\geq c$, we have
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]$ $\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]-c\gamma\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}+\varepsilon_{n}.$
Together with (15), the assumption $\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}:S_{\theta}]\|\geq\kappa$ implies that
$\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2}\geq\kappa-\varepsilon_{n}/\underline{\gamma}^{2}.$
Thus, for sufficiently large $n$, we have $\kappa-\varepsilon_{n}/\underline{\gamma}^{2}>0$ so that
$D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\mathcal{P}_{\gamma}[\theta]}\right]$
$\leq D_{\alpha}^{(n)}\left[\widetilde{I}_{n}^{z,\omega^{*}}\,:\,S_{\theta}\right]-\left(c-\frac{\varepsilon_{n}}{\underline{\gamma}(\kappa-\varepsilon_{n}/\underline{\gamma}^{2})}\right)\gamma\|\nabla_{\theta}D_{\alpha}^{(n)}[\widetilde{I}_{n}^{z,\omega^{*}}:S_{\theta}]\|^{2},$
which concludes the proof. ∎
Appendix D Proof of Proposition 3
The difference between $\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]$ and $\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]$ is explicitly written as
$\|\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]\|$ $=\|\{\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\}-\{\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]-\theta^{(0)}\}\|$
$=\frac{\gamma_{1}z}{n}\left\|\frac{\nabla_{\theta^{(0)}}\log S_{\theta}(\omega^{*})}{S_{\theta^{(0)}}(\omega^{*})}\right\|,$
which proves the first assertion. For the second assertion, observe
$\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg|_{\theta=\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}$
$=\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg|_{\theta=\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}-z\frac{\nabla_{\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega^{*})}{S_{\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega^{*})}.$
This gives the second assertion, which completes the proof. ∎
Appendix E Additional comparison of optimization paths
This appendix provides an additional comparison of optimization paths. Here we consider contaminated observations defined as
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,n=500$
with $(X_{t})$ generated from AR(2) model having $(\sigma,\varphi_{1},\varphi_{2})$.
Figure 7: Comparison of optimization paths using the AR(2) model with the length $n$ of a time series set to $500$. Panels (a) and (c) show the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $10000$ iterations (the inverted triangles), where the colors indicate different values of the contamination $z$. Panels (b) and (d) show the corresponding results for the Itakura–Saito divergence minimization.
To remove the parameter constraints arising from the stationarity condition, we employ the following transformation:
$\theta_{2}:=\,\mathrm{atanh}\left(\frac{\varphi_{1}}{1-\varphi_{2}}\right)\in\mathbb{R}\quad\text{and}\quad\theta_{3}:=\,\mathrm{atanh}(\varphi_{2})\in\mathbb{R}.$ (16)
This transformation is interpreted as follows. First, we transform the AR coefficients $\varphi_{1},\varphi_{2}$ into the partial autocorrelation (PARCOR) coefficients $\kappa_{1},\kappa_{2}$ using the Durbin–Levinson formula:
$\kappa_{1}:=\frac{\varphi_{1}}{1-\varphi_{2}}\in(-1,1)\quad\text{and}\quad\kappa_{2}:=\varphi_{2}\in(-1,1).$
Next, we map the PARCOR coefficients to unrestricted real values via the Fisher transformation:
$\theta_{2}:=\,\mathrm{atanh}\left(\kappa_{1}\right)\in\mathbb{R}\quad\text{and}\quad\theta_{3}:=\,\mathrm{atanh}(\kappa_{2})\in\mathbb{R},$
which yields the transformation in (16). We perform 10000 iterations of gradient descent with a step size of 0.01 with respect to $(\log\sigma,\theta_{2},\theta_{3})$.
Figure 7 shows a comparison of the optimization paths for the AR(2) model, where the true parameter values are set to $(\sigma^{*},\varphi^{*}_{1},\varphi^{*}_{2})=(1,0.9,-0.2)$ and the hyperparameter $\alpha$ is set to $\alpha=0.9$. Even for the PARCOR coefficients (in addition to $\sigma$), the optimization trajectory based on the Itakura–Saito divergence remains sensitive to outliers, whereas that based on the spectral Rényi divergence remains stable with respect to all parameters.
Appendix F Additional simulation studies
This appendix presents additional numerical experiments.
Tables 5 and 6 show the estimation results for the Brune spectral model with attenuation in Section 4.2 on the basis of the different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$. For almost all the settings of initial values, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence performs the best.
Table 5: The mean values of biases with standard deviations without any trend. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init. Values greater than 10 are denoted by *.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R($\alpha=0.50$) $\theta^{(0),4}$ -0.03 ($\pm$0.05) -0.11 ($\pm$0.07) 0.26 ($\pm$0.15)
R ($\alpha=0.75$) $\theta^{(0),4}$ 0.04 ($\pm$0.06) -0.30 ($\pm$0.03) 1.33 ($\pm$0.17)
R ($\alpha=0.90$) $\theta^{(0),4}$ 0.09 ($\pm$0.08) -0.39 ($\pm$0.03) 6.13 ($\pm$0.04)
IS with $I_{n}$ $\theta^{(0),4}$ * * *
IS with $I_{n}^{S}$ $\theta^{(0),4}$ * * *
R ($\alpha=0.50$) $\theta^{(0),5}$ -0.01($\pm$0.06) -0.18 ($\pm$0.05) 0.47 ($\pm$0.16)
R ($\alpha=0.75$) $\theta^{(0),5}$ 0.01 ($\pm$ 0.06) -0.21 ($\pm$0.05) 0.68 ($\pm$0.22)
R ($\alpha=0.90$) $\theta^{(0),5}$ 0.02 ($\pm$0.06) -0.24 ($\pm$0.04) 1.18 ($\pm$0.34)
IS with $I_{n}$ $\theta^{(0),5}$ 0.07 ($\pm$0.11) -0.08 ($\pm$0.12) 0.12 ($\pm$0.18)
IS with $I_{n}^{S}$ $\theta^{(0),5}$ -0.28 ($\pm$0.16) 1.84 ($\pm$1.66) 1.90 ($\pm$0.63)
Table 6: The mean values of biases with standard deviations with the trigonometric trends. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init. Values greater than 10 are denoted by *.
Init $\widehat{\sigma}-\sigma^{*}$ $\widehat{\omega_{c}}-\omega_{c}^{*}$ $\widehat{Q}-Q^{*}$
R ($\alpha=0.50$) $\theta^{(0),4}$ 0.21 ($\pm$0.07) -0.21 ($\pm$0.06) 0.24 ($\pm$0.14)
R ($\alpha=0.75$) $\theta^{(0),4}$ 0.61 ($\pm$0.09) -0.43 ($\pm$0.02) 1.14 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),4}$ 1.65 ($\pm$0.12) -0.60 ($\pm$0.02) 6.05 ($\pm$0.05)
IS with $I_{n}$ $\theta^{(0),4}$ * * *
IS with $I_{n}^{S}$ $\theta^{(0),4}$ * * *
R ($\alpha=0.50$) $\theta^{(0),5}$ 0.23 ($\pm$0.07) -0.26 ($\pm$0.04) 0.41 ($\pm$0.16)
R ($\alpha=0.75$) $\theta^{(0),5}$ 0.55 ($\pm$0.08) -0.35 ($\pm$0.04) 0.55 ($\pm$0.20)
R ($\alpha=0.90$) $\theta^{(0),5}$ 1.50 ($\pm$0.10) -0.50 ($\pm$0.03) 0.85 ($\pm$0.35)
IS with $I_{n}$ $\theta^{(0),5}$ 3.91 ($\pm$0.07) -0.34 ($\pm$0.02) -0.40 ($\pm$0.02)
IS with $I_{n}^{S}$ $\theta^{(0),5}$ 2.85 ($\pm$0.95) -0.17 ($\pm$0.51) 2.26 ($\pm$0.93)
Figures 8 and 9 display an instance of a set of the estimated spectral densities. In each figure, the following representations are used:
  • The solid gray curve represents the periodogram $I_{n}$;
  • the solid black curve denotes the true spectral density $S^{\mathrm{BA}}_{\theta^{*}}$;
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged in;
  • the solid red, salmon pink, and green curves display the spectral density with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged in, respectively.
For any initial value and for any strength of trends, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence performs the best.
Figure 8: Spectral densities with the estimates plugged in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=2.5$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 9: Spectral densities with the estimates plugged in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=25$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.